### v1

# Rethinking the Value of Agent-Generated Tests for LLM-Based Software Engineering Agents

Abstract.
Large Language Model (LLM) code agents increasingly resolve repository-level issues by iteratively editing code, invoking tools, and validating candidate patches. In these workflows, agents often write tests on the fly, a paradigm adopted by many high-ranking agents on the SWE-bench leaderboard. However, we observe that GPT-5.2, which writes almost no new tests, can even achieve performance comparable to top-ranking agents. This raises the critical question: whether such tests meaningfully improve issue resolution or merely mimic human testing practices while consuming a substantial interaction budget.
To reveal the impact of agent-written tests, we present an empirical study that analyzes agent trajectories across six state-of-the-art LLMs on SWE-bench Verified. Our results show that while test writing is commonly adopted, but resolved and unresolved tasks within the same model exhibit similar test-writing frequencies Furthermore, these tests typically serve as observational feedback channels, where agents prefer value-revealing print statements significantly more than formal assertion-based checks. Based on these insights, we perform a controlled experiment by revising the prompts of four agents to either increase or reduce test writing. The results suggest that changes in the volume of agent-written tests do not significantly change final outcomes. Taken together, our study reveals that current test-writing practices may provide marginal utility in autonomous software engineering tasks.
Large Language Model, Agent-Written Tests, Agent Trajectory Analysis, Software Development Agent
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Code agents are increasingly used as effective paradigm for resolving software issues, where Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025b; Chen and Jiang, 2024) are integrated into scaffold of tools and interaction protocols to edit real repositories, invoke tools, and attempt to resolve issues end-to-end (Yang et al., 2024a; Örwall, 2024; Wang et al., 2024b; Gao et al., 2025b; Hong et al., 2024; Qian et al., 2024; Ehrlich et al., 2025; Xia et al., 2024). In this paper, a code agent denotes an LLM coupled with external tools and an iterative action–observation loop, and the scaffold refers to the surrounding tool interface and interaction protocol that specifies the agent’s allowed actions and feedback. Among the diverse skills required by code agents, testing plays a critical role, which expose regressions, validate hypotheses, and provide a feedback loop during patch development (Zhang et al., 2024; Xia et al., 2024; Liu et al., 2024; Mu et al., 2025).
Figure 1. Overview of our study design. RQ1 profiles emergent testing behaviours (test writing frequency, timing, and execution analysis). RQ2 characterizes the feedback signals encoded in agent-written tests (assertions vs. value-revealing prints) and the types of assertions. RQ3 applies prompt interventions to encourage or discourage writing tests, and measures both outcome impact and efficiency impact.
When operating on repository-level tasks, agents typically use tests as a primary validation interface, which come from two main sources. The first is the repository’s existing, human-written test suite, which reflects developer intent and established project conventions (Jimenez et al., 2024; Chen and Jiang, 2025). The second is agent-written tests—new test artifacts written by the agent during problem solving that were not present in the original codebase. In contrast to curated human-written tests, agent-written tests are written on the fly during issue resolution, and their reliability depends on the model’s understanding of the specification, domain knowledge, and the semantics of the target codebase. Agent-written tests can be beneficial by surfacing edge cases and providing actionable feedback for fault localization and patch refinement. However, they can also be harmful if they embed incorrect assumptions or oracles, diverting effort toward satisfying the test rather than resolving the target issue. Moreover, test generation and execution introduce non-trivial overhead—consuming API calls and tokens and increasing context footprint—which can reduce the remaining budget available for core debugging and patching (Kim et al., 2025). When the resulting signals are low-value, this overhead may dilute the agent’s focus and become net detrimental.
To better understand agent-written tests, we conduct a preliminary quantitative analysis of agent trajectories on SWE-bench Verified (OpenAI, 2024) using mini-SWE-agent (SWE-agent Team, 2024), where testing is optional and not enforced by any hard-coded procedure. We find that agent-written testing is prevalent for several strong models. For example, claude-opus-4.5 (ranked #1 in this setting, 74.4% resolution) generates at least one new test artifact in about 83% of tasks. Surprisingly, we observe a pronounced contrast for gpt-5.2: it achieves a comparable resolution rate (71.8%), only 2.6 percentage points below claude-opus-4.5, while generating near-zero new tests (only in 0.6% of tasks). This interesting observation inspires us to a core question: Do agent-written tests truly facilitate task resolution, or do models merely mimic a learned software development practice while the resulting tests contribute little to the final patch? If the latter holds, then the widespread creation and execution of agent-written tests may represent a significant resource waste, consuming significant interaction budget without meaningful gains in task success. Therefore, we argue that a systematic empirical study is needed to understand the role of agent-written tests in resolving software issues.
Prior work mostly evaluates and benchmarks LLM-generated tests under predefined testing objectives and fixed quality metrics (e.g., unit tests, assertions, or issue-reproducing tests), typically with respect to a fixed target program or snapshot of code under test (Lops et al., 2025; Zhang et al., 2025b; Mündler et al., 2024; Yuan et al., 2024; Wang et al., 2025, 2024a; Schäfer et al., 2023). However, in complex real-world GitHub issue resolution (Jimenez et al., 2024; Khandpur et al., 2025), the codebase and candidate patches evolve over time, and test writing and usage arise dynamically as self-directed behaviors rather than pre-specified evaluation objectives. Yet the intrinsic tendency of high-autonomy agents to write and use tests during such issue resolution—and the impact of such agent-written tests on resolution outcomes—has not been systematically studied. This motivates a closer empirical investigation of agent-written tests, guided by the following three research questions.
Research questions and overview.
Guided by the above gap, we conduct a systematic empirical study of agent-written tests in GitHub issue resolution. Figure 1 summarizes our study design, which decomposes the problem into three complementary research questions. RQ1 characterizes the agents’ testing behaviors when test writing is required by the prompt: whether agents write tests, when they introduce them, and how intensively they execute them. RQ2 shifts from behaviors to test content, investigating what feedback signals agent-written tests actually emit at execution time (assertions vs. value-revealing prints) and what types of assertions agents use. RQ3 evaluates impact: by revising prompts to encourage or discourage writing new tests, we measure whether changing test-writing behavior meaningfully alters task resolution outcomes and what efficiency costs (API calls and tokens) these changes incur.
Summary of findings.
Across models, agent-written testing is best understood as a model-dependent process style rather than a dependable driver of success. First, RQ1 shows that test writing is a widespread behavior among the studied models, while resolved and unresolved trajectories within a model exhibit broadly similar test-writing rates. When agent-written tests exist, unsuccessful trajectories tend to spread test writing slightly more over the run and execute tests more frequently. Second, RQ2 reveals that agent-written tests primarily function as an observational feedback channel: value-revealing prints consistently dominate assert-based checks, and assertion forms concentrate on local-property and exact-value checks, with relational/range-style constraints remaining rare. Third, RQ3 provides controlled evidence that large shifts in whether agents write tests translate into only small shifts in task outcomes for most tasks, while efficiency effects can be substantial—inducing tests can increase interaction overhead without improving resolution, whereas suppressing tests can materially reduce API calls and token usage with only modest success losses.
Contributions.
The contribution of this work can be summarized as:
  • A behavioral analysis of agent-written tests from code agents. We characterize the agent-written testing behaviors of base LLM agents, including whether they create new test artifacts, when such test creation occurs within a trajectory, and how these tests are executed. Our results show that test writing and execution intensity are largely model-dependent process styles and only weakly align with task success (e.g., some high-performing models resolve many tasks while writing almost no tests).
  • A feedback-signal analysis of agent-written tests with a four-category assertion categorization. We separate verification-oriented assertions from observational outputs and introduce a rule-based AST classifier that maps assertions into four assertion categories. We find that tests largely serve an observational role: value-revealing prints consistently outnumber assertions; assertion usage is dominated by local-property and exact-value checks, whereas relational/range-style constraints are uncommon.
  • A causal evaluation of agent-written tests on task resolution. Through controlled prompt interventions that either encourage or suppress writing new test files, we quantify the causal effects of agent-written tests on task success and interaction efficiency. We demonstrate that large flips in test-writing status translate into only small changes in resolution outcomes for most tasks, whereas efficiency effects can be substantial—inducing tests can increase token and interaction overhead without improving success, while suppressing tests yields large cost reductions with only modest success drops.
Paper organization.
Section 2 describes our study setting, data collection, and measurement procedures. Section 3 characterizes emergent agent-written testing behaviors as process events—whether agents write tests, when they appear, and how intensively they are executed (RQ1). Section 4 shifts to test content by analyzing what feedback signals agent-written tests produce at execution time (assertions vs. value-revealing prints) and what types of assertions they use (RQ2). Section 5 evaluates the outcome and efficiency impacts of agent-written tests via prompt interventions that induce or suppress test writing (RQ3). Section 6 discusses implications, limitations, and future directions. Section 7 reviews related work, and Section 8 concludes the paper.
2. Methodology
In this section, we introduce the methodology for this study, including the benchmark, studied agent and LLMs, the extraction of agent-written tests, and implementation details. Our study is guided by three research questions:
  • RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
  • RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
  • RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
2.1. Benchmark
We use SWE-bench Verified as a standardized source of real-world software-engineering trajectories in which agents autonomously write and use tests when solving a GitHub issue (Jimenez et al., 2024). SWE-bench Verified is a filtered subset of SWE-bench that contains 500 instances. Each benchmark instance provides a GitHub issue, a fixed repository snapshot, and the official evaluation harness. We analyze agent-written test artifacts within the observed trajectories (OpenAI, 2024).
2.2. Agent and its LLMs
While many recent LLM-based agents incorporate curated testing components, such as specialized validation modules, dedicated test-planning stages, or multi-agent coordination (Liu et al., 2024; Zhang et al., 2024; Ruan et al., 2024; Cognition Labs, 2024), these frameworks can confound a model’s intrinsic tendencies with scaffold-induced constraints. To better isolate base-model behavior, we adopt mini-SWE-agent (SWE-bench Team, 2024; SWE-agent Team, 2024). It provides a lightweight agent work loop restricted to a standard bash interface: the agent interacts with the repository solely through the bash tool, executing commands in a bash shell (e.g., running python) and using standard command-line utilities to inspect and modify files. The model can create executable Python test files on the fly and run them via bash as part of its workflow. Crucially, mini-SWE-agent does not provide additional testing-specific functions or dedicated testing tools (e.g., test planners or structured testing modules), so the testing decisions (whether, when, and how) are left to the model. Accordingly, any observed behaviors (e.g., creating or running test artifacts) can be interpreted as model-native ones.
We select a diverse set of strong LLMs to capture heterogeneous agent-written testing behaviors under mini-SWE-agent. Specifically, we reference the SWE-bench Bash Only leaderboard [^1] https://www.swebench.com/ as of the time of writing (2025-12-11) and identify the top-six model families (rather than top entries, since a family may appear with multiple variants on the leaderboard). For each family, we use its highest-ranked model as the representative: claude-opus-4.5 (Anthropic, 2025) (74.4%), gemini-3-pro-preview (Google Cloud, 2025) (74.2%), gpt-5.2 (OpenAI, 2025) (71.8%), kimi-k2-thinking (Moonshot AI, 2025) (63.4%), minimax-m2 (MiniMax, 2025) (61.0%), and deepseek-v3.2-reasoner (DeepSeek, 2025a) (60.0%).
2.3. Data Extraction of Agent-Written Tests
In our study, agent-written tests are test files that an agent writes using the bash tool during task resolution. We extract agent-written tests from task trajectories, which are time-ordered interaction logs recorded during issue resolution that include the agent’s intermediate reasoning, concrete actions (e.g., bash commands), and the resulting observations. To find test files written during a trajectory, we scan the logged bash actions for file-writing operations, most commonly here-doc writes such as cat <<’EOF’ > path/to/file.pyEOF. We then keep only files whose paths match common Python test naming patterns, including filenames that start with test_ or end with _test.py or tests.py.
2.4. Implementation Details
We run all experiments using the official mini-SWE-agent codebase. All tasks are executed on a Linux server (Ubuntu 22.04.5) with an AMD Ryzen Threadripper PRO 7975WX CPU (32 cores / 64 threads), 251 GiB RAM. For model inference, we access the four LLMs through a combination of official provider APIs and the OpenRouter API. For evaluation, we use the official SWE-bench sb-cli tool to score each submitted patch under the benchmark harness. Across all experiments reported in this paper, the total LLM API cost is approximately USD 1,600.
3. RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
Motivation.
In a high-autonomy setting where testing is optional, agents may or may not write tests during issue resolution. RQ1 establishes a descriptive baseline of these emergent testing behaviors—what tests agents write, when they introduce them, and how intensively they run them. This baseline (i) clarifies what ”testing” looks like in this setting and (ii) provides grounded behavioral variables for later research questions.
Experiment Design.
RQ1 uses resolved vs. unresolved trajectories as a comparative lens to characterize systematic differences in test-related behaviors. We emphasize that these outcome-stratified comparisons are not intended to establish causality regarding task success; rather, they serve as a diagnostic tool to surface consistent differences in testing practices between successful and unsuccessful problem-solving processes. RQ1 reports descriptive summaries of three complementary aspects of test-oriented behavior:
  • Frequency (RQ1.1): whether the agent writes tests, and how many.
  • Timing (RQ1.2): when test writing happens during issue resolution.
  • Execution (RQ1.3): how intensively tests are run, and their outcomes.
3.1. RQ1.1 Frequency: Do Agents Write Test Artifacts?
Goal and measurements.
We examine whether base LLMs write test artifacts under a light scaffold. For each task, we record (i) whether the agent writes at least one test artifact, and (ii) if so, how many distinct test artifacts it writes. We report results separately for resolved and unresolved tasks.
Table 1. Per-model test writing rate by execution outcome
Resolved Unresolved All
Model #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests
claude-opus-4-5 372 314 (84.4%) 3.33 128 101 (78.9%) 4.12 500 415 (83.0%) 3.52
gemini-3-pro-preview 371 235 (63.3%) 2.02 129 73 (56.6%) 2.16 500 308 (61.6%) 2.05
gpt-5.2 359 3 (0.8%) 1.00 141 0 (0.0%) 500 3 (0.6%) 1.00
kimi-k2-thinking 317 309 (97.5%) 3.48 183 178 (97.3%) 3.83 500 487 (97.4%) 3.61
minimax-m2 305 302 (99.0%) 4.82 195 191 (97.9%) 5.76 500 493 (98.6%) 5.19
deepseek-v3.2-reasoner 300 277 (92.3%) 3.55 200 169 (84.5%) 4.08 500 446 (89.2%) 3.75
Notes. “Tasks w/ tests (count, %)” reports the number (and percentage) of tasks that write at least one test artifact within each outcome split. “Mean #tests” reports the mean number of distinct test artifacts, computed only over tasks that write at least one test artifact.
Results.
Table 1 shows that writing tests is common for most models, but not for gpt-5.2. Some models write tests in almost every task (e.g., minimax-m2 and kimi-k2-thinking). In contrast, gpt-5.2 almost never writes tests (3/500 tasks). Within the same model, resolved and unresolved tasks usually have similar test-writing rates. When tests are written, unresolved tasks often write as many or more distinct test artifacts than resolved tasks. This may reflect that harder tasks trigger more trial-and-error.
RQ1.1 Test Writing: Key Pattern Test writing is widespread across models in our high-autonomy issue-resolution setting. Most models write tests in a majority of tasks (61.6–98.6%), and four of six do so in at least 83% of tasks; gpt-5.2 is a clear outlier with near-zero test writing (0.6%).
3.2. RQ1.2 Timing: When Are Tests Written During the Run?
Goal and measurements.
Beyond whether tests are written (RQ1.1), we examine when test writing happens during a task execution. Writing tests in a tight window may look like a short ”checking phase”, while writing tests throughout the task may look like iterative debugging. This subsection is descriptive and does not claim effectiveness. We analyze only tasks that write at least one test artifact, so the timing metrics are defined. Because gpt-5.2 writes tests in only 3 tasks (RQ1.1), we omit its per-model timing summaries in RQ1.2 to avoid unstable estimates. We also exclude it from later analyses that require tasks with test writing. We use three normalized positions within the task: the first test-writing position, the last test-writing position, and their span:
$t_{\text{first}}$ $=\frac{\min(S_{\text{write}})}{N_{\text{steps}}},\quad t_{\text{last}}=\frac{\max(S_{\text{write}})}{N_{\text{steps}}}$
$s_{\text{write}}$ $=t_{\text{last}}-t_{\text{first}}=\frac{\max(S_{\text{write}})-\min(S_{\text{write}})}{N_{\text{steps}}}$
Here, $S_{\text{write}}$ is the set of step indices where the agent writes test artifacts, and $N_{\text{steps}}$ is the total number of interaction steps in the task. $t_{\text{first}}$ and $t_{\text{last}}$ are normalized positions in $[0,1]$. Smaller values mean the agent writes tests earlier in the task; larger values mean later. The span $s_{\text{write}}\in[0,1]$ measures how spread out test writing is across the task. Larger values mean more dispersed test writing; smaller values mean a more concentrated window.
Table 2. Per-model timing of test writing events
Resolved Unresolved
Model #Tasks w/ tests First test-writing position Last test-writing position Test-writing span #Tasks w/ tests First test-writing position Last test-writing position Test-writing span
claude-opus-4-5 314 0.34 0.75 0.41 101 0.30 0.78 0.48
gemini-3-pro-preview 235 0.53 0.67 0.14 73 0.55 0.70 0.15
kimi-k2-thinking 309 0.40 0.82 0.42 178 0.40 0.82 0.42
minimax-m2 302 0.35 0.86 0.51 191 0.29 0.85 0.56
deepseek-v3.2-reasoner 277 0.43 0.80 0.37 169 0.40 0.80 0.40
All models 1440 0.40 0.78 0.38 712 0.37 0.80 0.43
Notes. Values are macro-over-tasks means. “First test-writing position” and “Last test-writing position” are the normalized positions of the first and last test-writing events relative to total_steps. “Test-writing span” is the normalized distance between the first and last test-writing positions.
Results.
Table 2 summarizes test-writing positions for tasks that write tests. Across all models, the average first test-writing position is 0.40 for resolved tasks and 0.37 for unresolved tasks. The average last test-writing position is 0.78 (resolved) and 0.80 (unresolved). Models differ in when they start writing tests. For example, gemini-3-pro-preview starts later (0.53–0.55), while minimax-m2 and claude-opus-4-5 start earlier (0.29–0.35). Most models finish test writing late in the task (last position around 0.75–0.86). Models also differ in how spread out test writing is. gemini-3-pro-preview has a short span (0.14–0.15). minimax-m2 has a wider span (0.51–0.56), and claude-opus-4-5 is also relatively wide (0.41–0.48). kimi-k2-thinking is almost identical between resolved and unresolved tasks (0.40–0.82; span 0.42). Overall, unresolved tasks have a slightly larger average span than resolved tasks (0.43 vs. 0.38).
RQ1.2 Test-Writing Timing: Key Pattern Test writing typically finishes late, but its start time and span are mainly model-dependent; unresolved tasks are only slightly more spread out (0.43 vs. 0.38).
3.3. RQ1.3 Execution: How Intensively Are Agent-Written Tests Executed, and With What Process Outcomes?
Goal and measurements.
RQ1.3 describes how agents execute tests after they have written them. We measure (i) how often tests are executed, (ii) how often they are rerun relative to the number of written test artifacts, and (iii) how often executions fail at the process level. We treat an execution as failed if it ends with a non-zero return code (and successful otherwise). This captures execution friction during interaction with the environment, not patch correctness.
For each task $t$, let $E_{t}$ be the number of test executions, $A_{t}$ the number of agent-written test artifacts, and $F_{t}$ the number of executions with non-zero return codes. We report three task-level metrics: ExecCount ($E_{t}$), test executions per task; ExecPerTest ($E_{t}/A_{t}$), executions per written test artifact (rerun intensity); and FailRate ($F_{t}/E_{t}$), the fraction of executions that fail. We report macro-over-tasks means for each metric.
Table 3. Task-level execution effort and process-level outcomes of agent-written tests
Resolved Unresolved
Model #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%) #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%)
claude-opus-4-5 314 4.87 1.50 11.97 101 6.27 1.68 11.14
gemini-3-pro-preview 235 2.71 1.51 8.53 73 2.79 1.40 7.08
kimi-k2-thinking 309 5.39 1.62 24.95 178 6.54 1.76 21.05
minimax-m2 302 7.19 1.55 24.11 191 9.70 2.09 24.10
deepseek-v3.2-reasoner 277 3.74 1.11 27.37 169 4.66 1.32 29.55
All models 1440 4.89 1.46 19.68 712 6.52 1.70 21.05
Notes. #Tasks: tasks with test writing. Mean ExecCount: test executions per task. Mean ExecPerTest: executions per agent-written test artifact. Mean FailRate: % executions with non-zero return codes (macro-over-tasks).
Results.
Table 3 summarizes execution effort and process-level outcomes for tasks that write tests. Across all models, unresolved tasks execute tests more often than resolved tasks (Mean ExecCount: 6.52 vs. 4.89). They also rerun tests more per written test artifact (Mean ExecPerTest: 1.70 vs. 1.46). FailRate is slightly higher for unresolved tasks (21.05% vs. 19.68%). Models differ strongly in execution intensity. gemini-3-pro-preview runs tests the least (ExecCount $\approx$ 2.7–2.8). minimax-m2 runs tests the most, especially for unresolved tasks (ExecCount 9.70; ExecPerTest 2.09). FailRate also varies by model. claude-opus-4-5 and gemini-3-pro-preview have lower FailRate (about 7–12%), while deepseek-v3.2-reasoner, kimi-k2-thinking, and minimax-m2 are higher (about 21–30%).
RQ1.3 Test Execution: Key Pattern For tasks that write tests, unresolved tasks run tests more often and rerun them more per written test artifact (ExecCount 6.52 vs. 4.89; ExecPerTest 1.70 vs. 1.46), while process-level execution failures vary mainly by model (FailRate $\sim$7%–30%).
3.4. Summary of RQ1.
RQ1 provides a descriptive baseline of agent-written testing behaviors in a high-autonomy setting. Test writing is common for most models, but extremely rare for gpt-5.2 (Table 1). For tasks that write tests, test writing usually ends late in the task, while the start position and the writing span vary mainly by model (Table 2). Unresolved tasks are slightly more spread out in test writing on average (span 0.43 vs. 0.38). Finally, among tasks that write tests, unresolved tasks execute tests more often and rerun them more per written test artifact (Table 3). Process-level execution failures also vary by model (about 7–30% FailRate). RQ1, however, only describes when and how often tests appear and run, not what feedback they provide. RQ2 therefore analyzes the feedback encoded in agent-written tests.
4. RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
Motivation.
In our high-autonomy setting where testing is optional, tests may play different roles depending on the feedback they emit when executed. RQ1 treats tests as events in the trajectory—whether agents write them, when they appear, and how often they are run. RQ2 shifts to the content of those tests: the feedback they produce during execution. We capture this feedback through two common signals in agent-written tests: assertions (which fail when conditions are violated) and value-revealing prints (which expose runtime values). This view clarifies what agents use tests for when resolving GitHub issues.
Experiment Design.
RQ2 conditions on tasks that write at least one test artifact and reports descriptive summaries of two aspects of test feedback:
  • Signal counts (RQ2.1): how many feedback statements appear in agent-written tests, split into assertions vs. value-revealing prints.
  • Assertion types (RQ2.2): what types of assertions appear in agent-written tests, using a four-type categorization.
4.1. RQ2.1 Task-level feedback signal amount: How much feedback do agent-written tests encode?
Goal and measurements.
Conditioning on tasks that contain agent-written test artifact, we quantify how many feedback statements are encoded in those artifacts. We distinguish two signal types: (i) verification signals ($A$), i.e., assert statements that specify explicit checks, and (ii) observational signals ($P$), i.e., value-revealing print statements that expose runtime values or computed expressions. To ensure $P$ (prints) reflects observational feedback, we exclude pure-literal prints that emit only fixed strings (e.g., print("here")) and count only prints that expose runtime values, expressions, or execution results (e.g., print(obj.attr)). For each task $t$ with test artifacts $\mathcal{A}_{t}$, and for signal type $S\in\{A,P\}$, let $n^{S}_{t,a}$ be the number of signal statements of type $S$ in artifact $a\in\mathcal{A}_{t}$. We define the task-level signal totals:
$N^{S}_{t}=\sum_{a\in\mathcal{A}_{t}}n^{S}_{t,a},\qquad N^{\mathrm{total}}_{t}=N^{A}_{t}+N^{P}_{t}.$
We report macro-over-tasks means of $N^{A}_{t}$ (assertion count), $N^{P}_{t}$ (value-revealing print count), and $N^{\mathrm{total}}_{t}$ (overall signal count) for each model, separately for resolved and unresolved tasks.
Table 4. Task-level feedback signal amount encoded in agent-written tests
Resolved Unresolved Model #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) claude-opus-4-5 314 5.16 25.00 30.16 101 5.36 25.61 30.97 gemini-3-pro-preview 235 1.45 4.34 5.79 73 1.62 5.04 6.66 kimi-k2-thinking 309 2.86 20.72 23.57 178 3.51 24.03 27.54 minimax-m2 302 7.37 34.06 41.43 191 4.66 43.09 47.76 deepseek-v3.2-reasoner 277 3.51 16.43 19.94 169 3.31 20.95 24.27
Note. Macro-over-tasks means computed over tasks with tests. Assertions count assert statements; prints count value-revealing prints. $\bar{N}^{\mathrm{total}}=\bar{N}^{A}+\bar{N}^{P}$.
Figure 2. Composition of feedback signals in agent-written tests across models. Value-revealing prints dominate over assertions for all models.
Results.
Table 4 shows that, when agent-written tests are present, they can contain a substantial number of feedback statements per task (e.g., 19.94–47.76 total signals for several models). As shown in Figure 2, feedback is predominantly observational: for every model, value-revealing prints exceed assertions in the macro-average counts per task. Models differ markedly in overall signal volume. For example, minimax-m2 encodes the largest total signal counts (41.43–47.76), whereas gemini-3-pro-preview encodes far fewer (5.79–6.66), with other models in between. Across models, unresolved tasks tend to show slightly higher total signal counts, driven mainly by more value-revealing prints (e.g., deepseek-v3.2-reasoner: +4.33 total; minimax-m2: +6.33 total). Assertion counts are comparatively stable; notably, minimax-m2 has fewer assertions but more prints on unresolved tasks.
RQ2.1 Test Signal Amount: Key Takeaway When agents write tests, those tests often emit many runtime feedback statements per task, but the feedback is mostly observational: value-revealing prints consistently outnumber assertions across all models. Signal volume varies widely by model (e.g., minimax-m2 high vs. gemini-3-pro-preview low), and unresolved tasks show slightly more total signals mainly due to additional prints, while assertion counts remain relatively stable.
4.2. RQ2.2 Assertion categorization: What kinds of verification do assertions encode?
Goal and measurements.
RQ2.1 counts how many assert statements appear in agent-written tests, but counts alone do not tell us what those assertions check. Assertions can enforce different kinds of checks—for example, basic preconditions (e.g., non-None or type checks) versus checks against expected values or structures. Models may therefore differ not only in how often they assert, but also in what forms of checks they write. RQ2.2 provides a descriptive breakdown of assert statements into four assertion categories:
  • C1 Sanity checks. The assertion only checks existence or type, without constraining the expected behavior. Example: assert x is not None.
  • C2 Property checks. The assertion checks a property of a value or object (e.g., membership or validity) without fixing an exact output. Example: assert hasattr(obj, "attr").
  • C3 Relational checks. The assertion enforces a constraint such as a range, bound, or relationship between values. Example: assert 0 <= score <= 1. This category also includes checks that expect a specific exception, because they constrain the allowed behavior to “must fail with an exception of type EE” rather than matching a single concrete output.
  • C4 Exact checks. The assertion checks an exact value or deep structural equality. Example: assert output == expected_output.
To identify and categorize assertions, we implement a rule-based classifier over Python ASTs and map each extracted assertion to exactly one category. The classifier covers both native assert statements (e.g., assert a == b) and framework-provided assertion calls (e.g., self.assertEqual(a, b) in unittest). Concretely, for each test artifact, we parse the code into an AST and extract assertion events from: (i) native assert <expr> statements, and (ii) calls to framework assertion APIs. Some assert statements contain multiple checks in one line, combined with boolean operators (e.g., assert a > 0 and b == 1). In this example, a > 0 is a constraint check (C3) and b == 1 is an exact check (C4). For such compound assert statements, we decompose the expression into atomic checks and assign a single category by taking the highest category present under the ordering from C1 to C4, because the most specific check in the statement best reflects what the assertion is trying to enforce. Thus, assert a > 0 and b == 1 is labeled as C4.
Table 5. Assertion category distribution by model. Counts and percentages are computed over all assertion statements written by each model.
Model #Assertions C1 Sanity C2 Property C3 Relational C4 Exact # % # % # % # % claude-opus-4-5 2160 351 16.25% 807 37.36% 93 4.31% 909 42.08% gemini-3-pro-preview 458 76 16.59% 154 33.62% 36 7.86% 192 41.92% kimi-k2-thinking 1508 225 14.92% 622 41.25% 45 2.98% 616 40.85% minimax-m2 3117 618 19.83% 1291 41.42% 132 4.23% 1076 34.52% deepseek-v3.2-reasoner 1531 285 18.62% 537 35.08% 52 3.40% 657 42.91%
Note. C1–C4 denote four assertion categories defined by the form of the check (sanity, property, relational/approximate, and exact-output). Percentages are computed within each model, relative to the model’s total assertion count.
Results.
Table 5 shows that models have broadly similar assertion-category distributions. Across all five models, most assertions fall into C2 Property and C4 Exact, while C3 Relational remains consistently uncommon. For four models (claude-opus-4-5, gemini-3-pro-preview, kimi-k2-thinking, and deepseek-v3.2-reasoner), C4 Exact accounts for roughly 41–43% of assertions (40.85–42.91%), and C2 Property accounts for roughly 34–41% (33.62–41.25%). minimax-m2 follows the same overall shape but allocates a smaller share to C4 Exact (34.52%) and larger shares to C1 Sanity (19.83%) and C2 Property (41.42%). Across all models, C3 Relational is rare (2.98–7.86%), with the highest proportion in gemini-3-pro-preview (7.86%). Overall, the consistent scarcity of C3 may have several practical explanations. First, relational or approximate checks can be more delicate to specify and maintain than local property checks or direct equality checks. Second, such checks may be less common in everyday unit-test patterns that models imitate during code generation. Third, for many SWE-bench issues, agents may find it more straightforward to write either local property checks (C2) or exact expected outputs (C4) once they have a candidate fix. We treat these distributions as descriptive of which assertion forms appear in agent-written tests, not as evidence of correctness or impact on task resolution.
RQ2.2 Assertion Categorization: Key Takeaway Across models, assertions in agent-written tests are dominated by two forms: checks of local properties (e.g., checking that a required attribute is present) and checks against exact expected values. Assertions that express relational or range-style constraints are consistently uncommon (only a small single-digit fraction of assertions across models). minimax-m2 has a lower share of exact-value checks and a higher share of sanity-guard and local-property checks than the other models, but the overall pattern is similar across models.
4.3. Summary of RQ2
RQ2 shifts from when agents test (RQ1) to what feedback their agent-written tests produce at execution time. First, when tests are present, they often emit many feedback statements per task, but the feedback is dominated by value-revealing prints rather than assertions: across all models, print signals consistently outnumber assert checks, and total signal volume varies widely by model (RQ2.1). Second, when agents do write assertions, models show broadly similar mixes of assertion forms: most assertions either check local properties (e.g., the presence or validity of an attribute or field) or check exact expected values, while relational or range-style constraints are rare (RQ2.2). Taken together, RQ2 characterizes agent-written tests primarily as an observational feedback channel: most test feedback comes from value-revealing prints. This naturally raises the next question: do these agent-written tests meaningfully affect task resolution? We address this in RQ3.
5. RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
Motivation.
In RQ1, we find a weak alignment between agent-written tests and the final task success in this high-autonomy setting. For example, gpt-5.2 almost never writes new test artifacts (3/500 tasks, 0.6%), yet it still resolves 71.8% of tasks. In contrast, claude-opus-4.5 writes at least one new test artifact in about 83% of tasks, but its resolution rate is only 2.6 percentage points higher (74.4%). RQ2 further shows that, when tests are written, most feedback comes from value-revealing prints rather than assert-based checks. These findings raise a direct question: Do writing tests truly affect task resolution outcomes, and at what cost?
Experiment Design.
RQ3 answers two questions:
  • RQ3.1 (Outcome impact): If we encourage or discourage agents to write tests, how do task resolution outcomes change?
  • RQ3.2 (Efficiency impact): If we encourage or discourage agents to write tests, how do API calls and token usage change?
Model selection.
To isolate the effect of agent-written tests, we design two complementary intervention experiments: (i) encouraging agents to write tests, and (ii) discouraging agents from writing new test files. We choose models for each setup based on their baseline test-writing rate observed in RQ1 (Table 1), defined as the fraction of tasks where the agent writes test artifacts.
For the encourage test writing setup, we focus on low test-writing models and medium test-writing models, so there is meaningful headroom for increasing test creation. Specifically, we include gpt-5.2 (0.6%), an extreme low test-writing model in RQ1 with near-zero test creation. We also include gemini-3-pro-preview (61.1%), a medium test-writing model whose baseline test creation is already substantial but still leaves room for further increase.
For the discourage test writing setup, we start from high test-writing models that write tests in the vast majority of tasks in RQ1: four models show consistently high test-writing rates (83.0%–98.6%; Table 1). Due to budget constraints, we select two representatives from this group: kimi-k2-thinking (97.4%) and deepseek-v3.2-reasoner (89.2%).
Concretely, we use two prompt variants:
  • Encourage writing tests: for gpt-5.2 and gemini-3-pro-preview, we append an instruction to write at least one runnable new test file (a file whose name starts with test_ or ends with _test.py), separate from the repository’s existing tests.
  • Discourage writing tests: for kimi-k2-thinking and deepseek-v3.2-reasoner, we (i) remove the sentence “Test edge cases to ensure your fix is robust” and (ii) append an instruction to not write any new test files or scripts; robustness and edge cases should be handled using reasoning and code inspection only.
By comparing the performance of each agent under these revised prompts against their original performance, we isolate the specific impact of test writing. Consequently, the observed changes in behaviors or outcomes can be directly attributed to the agent-written tests.
5.1. RQ3.1 Does encouraging or discouraging test writing change task resolution?
Goal and measurements.
To answer whether encouraging or discouraging test writing changes task resolution, for each model, we compare each task under two conditions: the baseline run (under standard mini-SWE-agent prompt) and the intervention run (under our revised prompt). Specifically, we record two features: whether the run creates at least one new test artifact (No test vs. Has test) and whether the patch successfully resolve the issue (Fail vs. Success). Then, we analyze how these two features change from the baseline run to the intervention run. This respectively results in four possible transition groups for both test writing (No $\rightarrowNo test$, No $\rightarrowHas test$, Has $\rightarrowNo test$, Has $\rightarrowHas test$) and task outcome (Fail $\rightarrow$ Success, Success $\rightarrow$ Fail, Stable Success, and Stable Fail). To visualize the relationship between these shifts, we represent the results using a transition matrix. In this matrix, the rows represent the change in test-writing behavior, while the columns represent the change in task outcomes. This structure allows us to pinpoint the exact impact of our intervention. For instance, the intersection of No test $\rightarrow$ Has test and Fail $\rightarrow$ Success represents the instances where encouraging test-writing directly led to a breakthrough in resolving the task.
Table 6. Test-writing status flips and outcome transitions
[…]
Figure 3. Outcome-transition distribution on tasks with an intended test-status change
Results.
Our prompt interventions substantially change whether models write test artifacts, but these shifts rarely translate into outcome changes. As shown in Table 6, the encourage test writing prompt flips test status for the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview: 64.4% and 37.0% of tasks transition from No test to Has test, respectively. Conversely, the discourage test writing prompt removes tests at scale for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, moving 68.4% and 75.2% of tasks from Has test to No test.
Despite these large test-status shifts, resolution outcomes are largely stable. Figure 3 shows that, across models, an average of 83.2% of tasks keep the same final resolution result after the intervention. Table 6 further indicates that even when test status flips, success rates change only slightly. For example, for deepseek-v3.2-reasoner, discouraging test writing removes tests in 376 tasks but yields a net decrease of only 20 resolved tasks, a small change relative to the behavioral shift. Overall, changing how often a model writes test artifacts appears to be a weak lever for shifting task outcomes in this setting.
RQ3.1: Outcomes vs. Test-Status Changes Discouraging test writing removes agent-written tests at scale (at least 68% of tasks for kimi-k2-thinking and deepseek-v3.2-reasoner). Yet these large behavioral shifts rarely translate into outcome shifts: across models, 83.2% of tasks keep the same final resolution result after the intervention. Even when tests are induced, success does not increase correspondingly, suggesting that simply writing more test artifacts provides little leverage for fixing the underlying issue. Conversely, suppressing tests for hundreds of tasks does not cause outcomes to collapse; for deepseek-v3.2-reasoner, removing tests in 376 tasks yields only 20 fewer successes.
5.2. RQ3.2 How do API calls and token usage change?
Goal and measurements.
We further analyze the following three metrics based on the trajectories generated in RQ3.1: (i) average API calls per task, (ii) average input tokens per task, and (iii) average output tokens per task.
Table 7. API calls and token usage under baseline vs. encourage-tests / discourage-tests conditions.
Model Condition Tasks resolved Avg API Calls Avg Input Tokens Avg Output Tokens gpt-5.2 Baseline 359 (71.8%) 19.76 242,855 24,550 Encourage writing tests 359 (71.8%) 20.84 264,762 29,415 Change +0 (+0.0%) +1.08 (+5.5%) +21,907 (+9.0%) +4,866 (+19.8%) gemini-3-pro-preview Baseline 371 (74.2%) 40.33 666,096 11,114 Encourage writing tests 366 (73.2%) 39.21 641,307 10,943 Change -5 (-1.0%) -1.11 (-2.8%) -24,789 (-3.7%) -171 (-1.5%) kimi-k2-thinking Baseline 317 (63.4%) 46.82 668,449 14,895 Discourage writing tests 304 (60.8%) 30.25 340,689 8,468 Change -13 (-2.6%) -16.57 (-35.4%) -327,760 (-49.0%) -6,427 (-43.1%) deepseek-v3.2-reasoner Baseline 300 (60.0%) 46.40 637,297 52,120 Discourage writing tests 291 (58.2%) 35.06 427,780 44,823 Change -9 (-1.8%) -11.35 (-24.5%) -209,518 (-32.9%) -7,297 (-14.0%)
Note. Changes are computed as (Condition - Baseline). Encourage writing tests is applied to gpt-5.2 and gemini-3-pro-preview; Discourage writing tests is applied to kimi-k2-thinking and deepseek-v3.2-reasoner.
Results.
Table 7 reports task outcomes (Tasks resolved) and efficiency metrics (API calls and tokens) under the baseline and the two intervention conditions. Overall, the interventions have only marginal impact on resolution rates, but they can noticeably reshape efficiency—in ways that depend on each model’s baseline test-writing propensity.
For the encourage test writing setup, the effects differ between the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview. For gpt-5.2, encouraging test writing increases overhead: API calls rise by 5.5% and output tokens by 19.8%, while the resolution rate remains unchanged from baseline. In the medium test-writing model gemini-3-pro-preview, we observe very small shifts in overall usage (e.g., $1.5\%$ completion tokens). A likely reason is that this model already writes tests in about 61.6% of baseline tasks, leaving limited headroom for the prompt to flip tasks from No test to Has test; accordingly, the net change in testing-related interaction is small on average. Moreover, because the model is nondeterministic, even when a task contains tests in both baseline and encouraged runs, the amount of testing (and related interaction) can differ slightly between the paired runs; these small within-task fluctuations average out, keeping aggregate cost differences modest and making a small decrease reasonable.
The most striking efficiency shifts appear in the discourage test writing setup for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner. Discouraging test writing yields large reductions in resource consumption, with input tokens dropping by 49.0% and 32.9%, respectively. For kimi-k2-thinking, average API calls fall by 35.4%, cutting interaction workload by over a third. Notably, these efficiency gains come with only small decreases in success rates (2.6% for kimi-k2-thinking and 1.8% for deepseek-v3.2-reasoner), suggesting that a substantial portion of the baseline interaction budget spent on test writing for these models has limited marginal benefit for task resolution in this setting.
RQ3.2: Cost changes are larger than outcome changes Changing whether agents write tests has little effect on task resolution, but it can noticeably change efficiency. More tests increases overhead for low test-writing model models without improving outcomes: for gpt-5.2, API calls rise by 5.5% and output tokens by 19.8% while the resolution rate stays unchanged. In contrast, suppressing tests yields large efficiency gains for high test-writing models: input tokens drop by 49.0% (kimi-k2-thinking) and 32.9% (deepseek-v3.2-reasoner), and kimi-k2-thinking reduces API calls by 35.4%. These savings come with only small success drops (2.6% and 1.8%), implying that most test-writing overhead in these models delivers low marginal utility for producing a successful final patch.
Summary of RQ3.
Overall, more agent-written tests do not mean more solves in this high-autonomy setting. When we push models to write more tests, test-writing status flips at scale (e.g., 64.4% for gpt-5.2), yet task outcomes remain largely unchanged: on average, 83.2% of tasks keep the same success/fail result after the intervention. At the same time, more tests can be expensive: for the low test-writing model gpt-5.2, inducing test writing increases overhead (+5.5% API calls; +19.8% output tokens) without any gain in resolution. Conversely, fewer tests can be much cheaper with only small outcome losses: for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, suppressing test writing cuts input tokens by 49.0% and 32.9%, respectively, while success drops remain modest (2.6% and 1.8%). Taken together, varying the amount of agent-written tests strongly reshapes resource usage but has limited leverage on whether the final patch resolves the issue, suggesting that much of the test-writing effort provides low marginal utility for task resolution.
6. Discussion and Future Work
6.1. Implications
Our study indicates that the testing behaviors conducted by current LLMs, without clear guidance, do not show significant improvement in agent-based software issue resolution. It suggests two main implications for the research and practice in designing and utilizing LLM software agents.
On the one hand, our results reveal inherent limitations in how existing LLMs leverage testing for issue resolution when explicit guidance is absent. This highlights the importance of instructing agents not only on whether tests should be written, but also on how they should be designed and, more critically, how to effectively interpret and utilize the feedback generated by test execution.
On the other hand, given that current LLMs often fail to extract meaningful value from test writing, practitioners may consider adopting a more conservative approach to agent-generated tests. Such an approach could involve more sophisticated prompting strategies or guardrail mechanisms that help agents determine when test writing is truly necessary. Additionally, code agents could benefit from incorporating a cost–benefit monitoring framework that tracks the overhead associated with test-related activities, such as test writing, execution, and failure analysis, to assess their actual contribution to patch refinement.
6.2. Future work
Our findings motivate two future-work topics on agent-written testing in lightly scaffolded, high-autonomy setups.
Evaluating on-the-fly test quality in a non-stationary code state. Traditional test-quality metrics (e.g., coverage, mutation score, fault revelation) assume a fixed snapshot of the system under test and reproducible executions (Yu et al., 2023; Ryan et al., 2024; Shin et al., 2024; Bhatia et al., 2024; Yang et al., 2024b; Molinelli et al., 2025; Harman et al., 2025). In agentic development, the code state is non-stationary: tests are written and run against intermediate repository versions that may later be overwritten, and the final patch may not preserve the exact state a test originally validated. This complicates attribution (which version/function a test exercised), reproducibility, and the use of conventional quality pipelines. Future work should develop instrumentation for execution-time evaluation—e.g., capturing runnable snapshots (code, environment, commands) at test time—and define metrics that remain meaningful for transient intermediate artifacts.
Self-evolving test-generation strategies without over-constraining exploration. Our results highlight a tension between structure and autonomy: fully self-directed testing can increase cost with limited outcome gains, while a heavy, fixed workflow can narrow the search space and preclude better strategies. The goal is to produce higher-value tests by deciding when to test, what to verify (oracle design), and how to allocate budget across reasoning, editing, and validation, while preserving task-specific flexibility. A promising direction is self-evolving (Robeyns et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Hu et al., 2025) testing policies: instead of a static hand-written prompt, allow the agent to revise its own testing prompt/policy from environment feedback and failure modes, adapting to task context and model capabilities (Gao et al., 2025a). Future work can formalize this as closed-loop optimization under cost and safety constraints, and compare human-specified versus self-adapted strategies under matched budgets and controlled scaffolds.
6.3. Threats to Validity
Internal validity. Agent runs can vary due to stochastic decoding and tool/environment nondeterminism, which may affect both testing behavior and resolution outcomes. Moreover, success–failure comparisons can reflect differences in task difficulty or interaction length (e.g., debugging duration), not only testing. We mitigate these concerns by treating observational results as descriptive, by using prompt-only interventions that change test creation while keeping the agent setup fixed, and by reporting task-level outcome transitions (not only aggregate resolution deltas) to show where outcomes change versus remain stable.
External validity. Our findings are based on SWE-bench under a light scaffold and a specific set of models and providers; absolute magnitudes may differ under other benchmarks, programming languages, toolchains (e.g., enforced CI), or future model versions. To support transfer, we focus on patterns that may recur in similar agent settings (e.g., large cross-model differences in testing style, observation-dominant feedback, and limited outcome sensitivity under sizable shifts in test creation), and we provide precise measurement definitions and intervention prompts to facilitate replication under alternative setups.
Data construction validity. Measurements rely on explicit operational definitions and automated extraction. Test adoption is detected through newly created test-like files, and feedback signals are extracted via deterministic AST-based rules for assertions and value-bearing prints, including common helper-style assertion APIs and a tiered taxonomy for assertion forms. These procedures can miss unconventional test artifacts, project-specific helpers, or edge-case syntax patterns. We reduce construction error by using AST parsing (rather than surface regex heuristics), conservative rules for counting value-bearing prints, and fully deterministic, reproducible extraction and categorization; nevertheless, results should be interpreted with respect to these explicit definitions.
7. Related Work
Evaluation for LLM-Generated Tests. Prior work evaluates LLM-generated testing artifacts under predefined objectives, most commonly unit tests and assertions, via systems and empirical studies on test-suite quality and model/prompt improvements (Lops et al., 2025; Yuan et al., 2024; Schäfer et al., 2023; Li et al., 2025; Yang et al., 2025), including targeted oracle generation such as assertions (Zhang et al., 2025b). Recent surveys further systematize this space, summarizing how requirements artifacts are translated into tests and the quality criteria used to judge generated tests (Yang et al., 2025). Complementing academic evaluations, industrial studies report closed-loop pipelines that combine LLM-based test generation with mutation-guided feedback to steer or refine generated tests toward stronger fault-revealing capability (Harman et al., 2025). These studies typically score outputs with fixed quality metrics (e.g., coverage, mutation-based adequacy proxies, fault revelation) on a fixed target program or code snapshot. Benchmarks similarly cast testing as a standalone objective with fixed tasks and protocols (e.g., TestEval (Wang et al., 2025), SWT-bench (Mündler et al., 2024)). In contrast, our study focuses on agent-written tests that emerge dynamically during high-autonomy, multi-step resolution of real-world GitHub issues, where the codebase and candidate patches evolve over time. We treat test writing and execution as an emergent process behavior, and characterize (i) whether/when/how intensively agents test, (ii) what signals tests encode (assertions vs. observational prints), and (iii) how these behaviors relate to resolution outcomes.
Trajectory Analysis of Software Agents. Recent work moves beyond final patch and binary success/failure by analyzing the intermediate reasoning and execution traces of LLM-based agents. Studies have examined action–observation patterns that distinguish successful from failed runs (Bouzenia and Pradel, 2025), compared trajectory length and fault-localization accuracy across agents (Majgaonkar et al., 2026), and proposed workflow taxonomies that decompose agent behavior into stages such as localization, patching, and testing-related steps (Ceka et al., 2025). Others conduct systematic failure analyses that identify root causes such as diagnostic errors and unproductive loops (Liu et al., 2025), while process-oriented studies further show that agents often hit recurrent execution errors during issue resolution, motivating lightweight checks and recovery components for robustness (Chen et al., 2026). Overall, existing trajectory analyses emphasize action sequences, outcome separation, and error categorization, but rarely examine whether and how agents autonomously decide to test. Our work addresses this gap by characterizing the emergent testing behaviors observed in execution trajectories and analyzing what feedback agent-written tests actually provide during issue resolution.
8. Conclusion
This paper revisited the common intuition that ”testing helps” for LLM-based software agents in a high-autonomy setting where writing and running tests is not specified in the prompt. Using SWE-bench Verified trajectories produced under a light agent scaffold, we separated (i) whether/when/how agents write and execute tests, (ii) what feedback those tests encode, and (iii) whether changing test-writing instructions actually shifts task outcomes and efficiency. RQ1 established a descriptive baseline of emergent testing behaviors. Across models, test writing propensity differs dramatically (from near-universal test writing to almost none), while within the same model, resolved and unresolved tasks show broadly similar test-writing rates. When tests are present, unresolved tasks tend to spread test writing across a slightly larger portion of the trajectory and execute tests more often, with higher re-execution intensity; process-level execution failures, however, vary mainly by model rather than outcome. These results highlight that in a high-autonomy setting, ”testing” is a model-specific process behavior and is not tightly coupled to eventual success. RQ2 then showed that the content of agent-written tests is largely observational. Across all models, value-revealing prints consistently outnumber assert-based checks, and overall signal volume varies widely by model. When assertions are used, their forms are broadly similar across models: most are local property checks or exact-value checks, while relational or range-style constraints are consistently rare. Taken together, agent-written tests function primarily as a runtime probing interface rather than a systematic verification mechanism. Finally, RQ3 provided controlled evidence about impact. Prompt-only interventions flip test-writing status at scale—inducing tests for the low test writing model and suppressing test writing for high test writing models—yet task outcomes are mostly unchanged (the majority of tasks preserve the same success/fail result). In contrast, efficiency effects can be substantial: for gpt-5.2—the near-zero test-writing model—encouraging test writing increases overhead without improving resolution, whereas for test-heavy models (e.g., kimi-k2-thinking and deepseek-v3.2-reasoner) discouraging test writing can sharply reduce API calls and token usage with only modest decreases in success. Overall, more agent-written tests do not mean more solves: varying test-writing effort primarily reshapes resource usage and interaction patterns, with limited leverage on whether the final patch resolves the issue.
Overall, our findings provide an empirical answer to the question: Help or Habit? In high-autonomy software issue resolution, agent-written tests often behave more like a reproduced software development lifecycle routine than a dependable source of help: models that write many tests are not consistently more successful, and prompt interventions that induce or suppress test writing rarely change whether the issue is resolved. What tests do change is the process footprint—API calls, token usage, and interaction steps—indicating that test writing frequently reflects how an agent chooses to work, not whether it can reliably validate the patch.
9. Data Availability
To support transparency and reproducibility, we publicly release a replication package containing the datasets, raw trajectories, and analysis scripts used in this study. The package is available on Figshare: https://figshare.com/s/8fa6412b86bf54fcd27e?file=61474897.

### v2

@@
 Large Language Model (LLM) code agents increasingly resolve repository-level issues by iteratively editing code, invoking tools, and validating candidate patches.
-In these workflows, agents often write tests on the fly, a paradigm adopted by many high-ranking agents on the SWE-bench leaderboard.
-However, we observe that GPT-5.2, which writes almost no new tests, can even achieve performance comparable to top-ranking agents.
-This raises the critical question: whether such tests meaningfully improve issue resolution or merely mimic human testing practices while consuming a substantial interaction budget.
-To reveal the impact of agent-written tests, we present an empirical study that analyzes agent trajectories across six state-of-the-art LLMs on SWE-bench Verified.
-Our results show that while test writing is commonly adopted, but resolved and unresolved tasks within the same model exhibit similar test-writing frequencies Furthermore, these tests typically serve as observational feedback channels, where agents prefer value-revealing print statements significantly more than formal assertion-based checks.
-Based on these insights, we perform a controlled experiment by revising the prompts of four agents to either increase or reduce test writing.
-The results suggest that changes in the volume of agent-written tests do not significantly change final outcomes.
-Taken together, our study reveals that current test-writing practices may provide marginal utility in autonomous software engineering tasks.
+In these workflows, agents often write tests on the fly, but the value of this behavior remains unclear.
+For example, GPT-5.2 writes almost no new tests yet achieves performance comparable to top-ranking agents.
+This raises a central question: do such tests meaningfully improve issue resolution, or do they mainly mimic a familiar software-development practice while consuming interaction budget?
+To better understand the role of agent-written tests, we analyze trajectories produced by six strong LLMs on SWE-bench Verified.
+Our results show that test writing is common, but resolved and unresolved tasks within the same model exhibit similar test-writing frequencies.
+When tests are written, they mainly serve as observational feedback channels, with value-revealing print statements appearing much more often than assertion-based checks.
+Based on these insights, we perform a prompt-intervention study by revising the prompts used with four models to either increase or reduce test writing.
+The results suggest that prompt-induced changes in the volume of agent-written tests do not significantly change final outcomes in this setting.
+Taken together, these results suggest that current agent-written testing practices reshape process and cost more than final task outcomes.
 Large Language Model, Agent-Written Tests, Agent Trajectory Analysis, Software Development Agent
-copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2026; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
+journalyear: 2026conference: ; October, 2026; Washington, DC, USA
 1. Introduction
-Code agents are increasingly used as effective paradigm for resolving software issues, where Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025b; Chen and Jiang, 2024) are integrated into scaffold of tools and interaction protocols to edit real repositories, invoke tools, and attempt to resolve issues end-to-end (Yang et al., 2024a; Örwall, 2024; Wang et al., 2024b; Gao et al., 2025b; Hong et al., 2024; Qian et al., 2024; Ehrlich et al., 2025; Xia et al., 2024).
+Code agents increasingly tackle software issues by combining Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025b; Chen and Jiang, 2024) with tools and interaction protocols that let them edit real repositories, invoke tools, and attempt end-to-end issue resolution (Yang et al., 2024a; Örwall, 2024; Wang et al., 2024b; Gao et al., 2025b; Hong et al., 2024; Qian et al., 2024; Ehrlich et al., 2025; Xia et al., 2024).
 In this paper, a code agent denotes an LLM coupled with external tools and an iterative action–observation loop, and the scaffold refers to the surrounding tool interface and interaction protocol that specifies the agent’s allowed actions and feedback.
-Among the diverse skills required by code agents, testing plays a critical role, which expose regressions, validate hypotheses, and provide a feedback loop during patch development (Zhang et al., 2024; Xia et al., 2024; Liu et al., 2024; Mu et al., 2025).
+Among the diverse skills required by code agents, testing plays a critical role: it exposes regressions, validates hypotheses, and provides a feedback loop during patch development (Zhang et al., 2024; Xia et al., 2024; Liu et al., 2024; Mu et al., 2025).
 Figure 1.
-Overview of our study design.
-RQ1 profiles emergent testing behaviours (test writing frequency, timing, and execution analysis).
-RQ2 characterizes the feedback signals encoded in agent-written tests (assertions vs.
-value-revealing prints) and the types of assertions.
-RQ3 applies prompt interventions to encourage or discourage writing tests, and measures both outcome impact and efficiency impact.
+Overview of the study design.
+RQ1 examines testing behaviors, RQ2 analyzes feedback signals in agent-written tests, and RQ3 studies observed outcome and efficiency changes under test-writing interventions.
 When operating on repository-level tasks, agents typically use tests as a primary validation interface, which come from two main sources.
@@
 When the resulting signals are low-value, this overhead may dilute the agent’s focus and become net detrimental.
-To better understand agent-written tests, we conduct a preliminary quantitative analysis of agent trajectories on SWE-bench Verified (OpenAI, 2024) using mini-SWE-agent (SWE-agent Team, 2024), where testing is optional and not enforced by any hard-coded procedure.
-We find that agent-written testing is prevalent for several strong models.
-For example, claude-opus-4.5 (ranked #1 in this setting, 74.4% resolution) generates at least one new test artifact in about 83% of tasks.
-Surprisingly, we observe a pronounced contrast for gpt-5.2: it achieves a comparable resolution rate (71.8%), only 2.6 percentage points below claude-opus-4.5, while generating near-zero new tests (only in 0.6% of tasks).
-This interesting observation inspires us to a core question: Do agent-written tests truly facilitate task resolution, or do models merely mimic a learned software development practice while the resulting tests contribute little to the final patch?
-If the latter holds, then the widespread creation and execution of agent-written tests may represent a significant resource waste, consuming significant interaction budget without meaningful gains in task success.
+To better understand agent-written tests, we conduct a quantitative analysis of agent trajectories on SWE-bench Verified (OpenAI, 2024) using mini-SWE-agent (SWE-agent Team, 2024), where testing is optional and not enforced by any hard-coded procedure.
+Agent-written testing is prevalent for several strong models.
+For example, Claude Opus 4.5 (ranked #1 in this setting, 74.4% resolution) generates at least one new test artifact in about 83% of tasks.
+By contrast, GPT-5.2 achieves a comparable resolution rate (71.8%), only 2.6 percentage points below Claude Opus 4.5, while generating near-zero new tests (only in 0.6% of tasks).
+This observation motivates a core question: when prompts and model tendencies lead agents to write more tests, do the observed outcomes actually improve, or do models merely mimic a learned software-development practice while the resulting tests contribute little to the final patch?
+If the latter holds, then the widespread creation and execution of agent-written tests may represent a substantial waste of resources, consuming interaction budget without meaningful gains in task success.
 Therefore, we argue that a systematic empirical study is needed to understand the role of agent-written tests in resolving software issues.
@@
 However, in complex real-world GitHub issue resolution (Jimenez et al., 2024; Khandpur et al., 2025), the codebase and candidate patches evolve over time, and test writing and usage arise dynamically as self-directed behaviors rather than pre-specified evaluation objectives.
-Yet the intrinsic tendency of high-autonomy agents to write and use tests during such issue resolution—and the impact of such agent-written tests on resolution outcomes—has not been systematically studied.
+Yet the intrinsic tendency of high-autonomy agents to write and use tests during such issue resolution—and how prompt-induced changes in test writing relate to observed resolution outcomes and costs—has not been systematically studied.
 This motivates a closer empirical investigation of agent-written tests, guided by the following three research questions.
 Research questions and overview.
-Guided by the above gap, we conduct a systematic empirical study of agent-written tests in GitHub issue resolution.
+Guided by this gap, we study the role of agent-written tests in GitHub issue resolution.
 Figure 1 summarizes our study design, which decomposes the problem into three complementary research questions.
-RQ1 characterizes the agents’ testing behaviors when test writing is required by the prompt: whether agents write tests, when they introduce them, and how intensively they execute them.
+RQ1 characterizes the agents’ testing behaviors under a light scaffold where test writing is optional: whether agents write tests, when they introduce them, and how intensively they execute them.
 RQ2 shifts from behaviors to test content, investigating what feedback signals agent-written tests actually emit at execution time (assertions vs.
-value-revealing prints) and what types of assertions agents use.
-RQ3 evaluates impact: by revising prompts to encourage or discourage writing new tests, we measure whether changing test-writing behavior meaningfully alters task resolution outcomes and what efficiency costs (API calls and tokens) these changes incur.
+value-revealing prints), what types of assertions they use, and what kinds of runtime information those value-revealing prints typically inspect.
+RQ3 examines observed outcome and efficiency changes: by revising prompts to encourage or discourage writing new tests, we measure how changing test-writing behavior is accompanied by changes in task resolution outcomes and efficiency costs (API calls and tokens).
 Summary of findings.
 Across models, agent-written testing is best understood as a model-dependent process style rather than a dependable driver of success.
-First, RQ1 shows that test writing is a widespread behavior among the studied models, while resolved and unresolved trajectories within a model exhibit broadly similar test-writing rates.
-When agent-written tests exist, unsuccessful trajectories tend to spread test writing slightly more over the run and execute tests more frequently.
-Second, RQ2 reveals that agent-written tests primarily function as an observational feedback channel: value-revealing prints consistently dominate assert-based checks, and assertion forms concentrate on local-property and exact-value checks, with relational/range-style constraints remaining rare.
-Third, RQ3 provides controlled evidence that large shifts in whether agents write tests translate into only small shifts in task outcomes for most tasks, while efficiency effects can be substantial—inducing tests can increase interaction overhead without improving resolution, whereas suppressing tests can materially reduce API calls and token usage with only modest success losses.
-Contributions.
-The contribution of this work can be summarized as:
+RQ1 shows that test writing is widespread but only weakly aligned with success.
+RQ2 shows that agent-written tests primarily serve as an observational feedback channel, with value-revealing prints dominating assert-based checks.
+RQ3 shows that prompt-induced changes in test writing have only small observed effects on task outcomes for most tasks, but can materially change efficiency.
+The contributions of this work can be summarized as follows:
   • A behavioral analysis of agent-written tests from code agents.
@@
 Our results show that test writing and execution intensity are largely model-dependent process styles and only weakly align with task success (e.g., some high-performing models resolve many tasks while writing almost no tests).
-  • A feedback-signal analysis of agent-written tests with a four-category assertion categorization.
-We separate verification-oriented assertions from observational outputs and introduce a rule-based AST classifier that maps assertions into four assertion categories.
-We find that tests largely serve an observational role: value-revealing prints consistently outnumber assertions; assertion usage is dominated by local-property and exact-value checks, whereas relational/range-style constraints are uncommon.
-  • A causal evaluation of agent-written tests on task resolution.
-Through controlled prompt interventions that either encourage or suppress writing new test files, we quantify the causal effects of agent-written tests on task success and interaction efficiency.
-We demonstrate that large flips in test-writing status translate into only small changes in resolution outcomes for most tasks, whereas efficiency effects can be substantial—inducing tests can increase token and interaction overhead without improving success, while suppressing tests yields large cost reductions with only modest success drops.
-Paper organization.
-Section 2 describes our study setting, data collection, and measurement procedures.
-Section 3 characterizes emergent agent-written testing behaviors as process events—whether agents write tests, when they appear, and how intensively they are executed (RQ1).
-Section 4 shifts to test content by analyzing what feedback signals agent-written tests produce at execution time (assertions vs.
-value-revealing prints) and what types of assertions they use (RQ2).
-Section 5 evaluates the outcome and efficiency impacts of agent-written tests via prompt interventions that induce or suppress test writing (RQ3).
-Section 6 discusses implications, limitations, and future directions.
-Section 7 reviews related work, and Section 8 concludes the paper.
+  • A feedback-signal analysis of agent-written tests covering assertions and value-revealing prints.
+We separate verification-oriented assertions from observational outputs and introduce rule-based AST analyses that map assertions into four categories and value-revealing prints into three coarse categories.
+We find that tests largely serve an observational role: value-revealing prints consistently outnumber assertions, those prints are dominated by value/content inspection and exception/status signals, and assertion usage is dominated by local-property and exact-value checks.
+  • A prompt-intervention study of agent-written tests on observed outcomes and efficiency.
+Through controlled prompt interventions that either encourage or suppress writing new test files, we study how changes in test-writing cues relate to observed task success and interaction efficiency under the same agent scaffold.
+We show that large flips in test-writing status are accompanied by only small observed changes in resolution outcomes for most tasks, whereas efficiency changes can be substantial.
+Inducing tests can increase token and interaction overhead without improving success, while suppressing tests yields large cost reductions with only modest success drops.
 2. Methodology
-In this section, we introduce the methodology for this study, including the benchmark, studied agent and LLMs, the extraction of agent-written tests, and implementation details.
-Our study is guided by three research questions:
-  • RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
-  • RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
-  • RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
+In this section, we introduce the methodology for this study, including the benchmark, the studied agent and LLMs, the extraction of agent-written tests, and implementation details.
+Our study is guided by three research questions: RQ1: What Testing Behaviors Emerge Under a Light Agent Scaffold?
+RQ2: What Feedback Signals Do Agent-Written Tests Provide?
+RQ3: How Does Prompting Test Writing Change Observed Outcomes and Costs?
 2.1. Benchmark
-We use SWE-bench Verified as a standardized source of real-world software-engineering trajectories in which agents autonomously write and use tests when solving a GitHub issue (Jimenez et al., 2024).
-SWE-bench Verified is a filtered subset of SWE-bench that contains 500 instances.
-Each benchmark instance provides a GitHub issue, a fixed repository snapshot, and the official evaluation harness.
-We analyze agent-written test artifacts within the observed trajectories (OpenAI, 2024).
-2.2. Agent and its LLMs
+We use SWE-bench Verified as our benchmark.
+The original SWE-bench benchmark is built from resolved GitHub issues drawn from 12 open-source Python repositories (Jimenez et al., 2024).
+SWE-bench Verified is a 500-instance subset released after a human-screening effort led by OpenAI in collaboration with the SWE-bench authors (OpenAI, 2024).
+Each instance provides a GitHub issue, a fixed repository snapshot, and the official evaluation harness.
+We analyze agent-written test artifacts within the resulting trajectories.
+$2.$ Agent and its LLMs
 While many recent LLM-based agents incorporate curated testing components, such as specialized validation modules, dedicated test-planning stages, or multi-agent coordination (Liu et al., 2024; Zhang et al., 2024; Ruan et al., 2024; Cognition Labs, 2024), these frameworks can confound a model’s intrinsic tendencies with scaffold-induced constraints.
@@
 The model can create executable Python test files on the fly and run them via bash as part of its workflow.
-Crucially, mini-SWE-agent does not provide additional testing-specific functions or dedicated testing tools (e.g., test planners or structured testing modules), so the testing decisions (whether, when, and how) are left to the model.
+Crucially, although the default mini-SWE-agent prompt includes a brief natural-language recommendation such as “Test edge cases to ensure your fix is robust,” this instruction is only advisory: it does not introduce any testing-specific function, dedicated testing tool, or hard-coded workflow component (e.g., a test planner, structured testing module, or enforced test-execution stage).
+Thus, in our setting, testing remains optional at runtime: the model may follow, delay, or ignore that recommendation, and the decisions of whether, when, and how to test are still left to the model.
 Accordingly, any observed behaviors (e.g., creating or running test artifacts) can be interpreted as model-native ones.
 We select a diverse set of strong LLMs to capture heterogeneous agent-written testing behaviors under mini-SWE-agent.
-Specifically, we reference the SWE-bench Bash Only leaderboard [^1] https://www.swebench.com/ as of the time of writing (2025-12-11) and identify the top-six model families (rather than top entries, since a family may appear with multiple variants on the leaderboard).
+For model selection, we use the SWE-bench Bash Only leaderboard with a cutoff date of 2025-12-11111https://www.swebench.com/ and identify the top-six model families (rather than top entries, since a family may appear with multiple variants on the leaderboard).
 For each family, we use its highest-ranked model as the representative: claude-opus-4.5 (Anthropic, 2025) (74.4%), gemini-3-pro-preview (Google Cloud, 2025) (74.2%), gpt-5.2 (OpenAI, 2025) (71.8%), kimi-k2-thinking (Moonshot AI, 2025) (63.4%), minimax-m2 (MiniMax, 2025) (61.0%), and deepseek-v3.2-reasoner (DeepSeek, 2025a) (60.0%).
+In the remainder of the paper, we refer to these models as Claude Opus 4.5, Gemini 3 Pro, GPT-5.2, Kimi K2 Thinking, MiniMax M2, and DeepSeek v3.2 Reasoner.
+In tables and figures, we further shorten Claude Opus 4.5, Kimi K2 Thinking, and DeepSeek v3.2 Reasoner to Claude 4.5, Kimi K2-T, and DeepSeek v3.2-R to save space.
+In each case, the shortened form still uniquely identifies the official model version evaluated in this study, without introducing ambiguity with another model release.
 2.3. Data Extraction of Agent-Written Tests
@@
 To find test files written during a trajectory, we scan the logged bash actions for file-writing operations, most commonly here-doc writes such as cat <<’EOF’ > path/to/file.pyEOF.
-We then keep only files whose paths match common Python test naming patterns, including filenames that start with test_ or end with _test.py or tests.py.
+We then keep only files whose paths match common Python test naming patterns, including filenames that start with test_ or end with _test.py or tests.py (pytest developers, 2026).
 2.4. Implementation Details
@@
 All tasks are executed on a Linux server (Ubuntu 22.04.5) with an AMD Ryzen Threadripper PRO 7975WX CPU (32 cores / 64 threads), 251 GiB RAM.
-For model inference, we access the four LLMs through a combination of official provider APIs and the OpenRouter API.
+For model inference, we access LLMs through a combination of official provider APIs and the OpenRouter API.
 For evaluation, we use the official SWE-bench sb-cli tool to score each submitted patch under the benchmark harness.
@@
 Resolved Unresolved All
-Model #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests #Tasks Tasks w/ tests (count, %) Mean #tests
-claude-opus-4-5 372 314 (84.4%) 3.33 128 101 (78.9%) 4.12 500 415 (83.0%) 3.52
-gemini-3-pro-preview 371 235 (63.3%) 2.02 129 73 (56.6%) 2.16 500 308 (61.6%) 2.05
+Model #Tasks Tasks w/ tests Mean #tests #Tasks Tasks w/ tests Mean #tests #Tasks Tasks w/ tests Mean #tests
+Claude 4.5 372 314 (84.4%) 3.33 128 101 (78.9%) 4.12 500 415 (83.0%) 3.52
+Gemini 3 Pro 371 235 (63.3%) 2.02 129 73 (56.6%) 2.16 500 308 (61.6%) 2.05
 GPT-5.2 359 3 (0.8%) 1.00 141 0 (0.0%) 500 3 (0.6%) 1.00
-kimi-k2-thinking 317 309 (97.5%) 3.48 183 178 (97.3%) 3.83 500 487 (97.4%) 3.61
-minimax-m2 305 302 (99.0%) 4.82 195 191 (97.9%) 5.76 500 493 (98.6%) 5.19
-deepseek-v3.2-reasoner 300 277 (92.3%) 3.55 200 169 (84.5%) 4.08 500 446 (89.2%) 3.75
-Notes.
-“Tasks w/ tests (count, %)” reports the number (and percentage) of tasks that write at least one test artifact within each outcome split.
-“Mean #tests” reports the mean number of distinct test artifacts, computed only over tasks that write at least one test artifact.
+Kimi K2-T 317 309 (97.5%) 3.48 183 178 (97.3%) 3.83 500 487 (97.4%) 3.61
+MiniMax M2 305 302 (99.0%) 4.82 195 191 (97.9%) 5.76 500 493 (98.6%) 5.19
+DeepSeek v3.2-R 300 277 (92.3%) 3.55 200 169 (84.5%) 4.08 500 446 (89.2%) 3.75
+Notes. “Tasks w/ tests” reports count and percentage within each outcome split. “Mean #tests” is computed only over tasks that write at least one test artifact.
 Results.
 Table 1 shows that writing tests is common for most models, but not for GPT-5.2.
-Some models write tests in almost every task (e.g., minimax-m2 and kimi-k2-thinking).
+Some models write tests in almost every task (e.g., MiniMax M2 and Kimi K2 Thinking).
 In contrast, GPT-5.2 almost never writes tests (3/500 tasks).
@@
 This may reflect that harder tasks trigger more trial-and-error.
-RQ1.1 Test Writing: Key Pattern Test writing is widespread across models in our high-autonomy issue-resolution setting.
-Most models write tests in a majority of tasks (61.6–98.6%), and four of six do so in at least 83% of tasks; gpt-5.2 is a clear outlier with near-zero test writing (0.6%).
+RQ1.1 Test Writing: Key Pattern Test writing is common for most models in this high-autonomy setting, but GPT-5.2 is a clear outlier with near-zero test writing.
 3.2. RQ1.2 Timing: When Are Tests Written During the Run?
@@
 Resolved Unresolved
-Model #Tasks w/ tests First test-writing position Last test-writing position Test-writing span #Tasks w/ tests First test-writing position Last test-writing position Test-writing span
-claude-opus-4-5 314 0.34 0.75 0.41 101 0.30 0.78 0.48
-gemini-3-pro-preview 235 0.53 0.67 0.14 73 0.55 0.70 0.15
-kimi-k2-thinking 309 0.40 0.82 0.42 178 0.40 0.82 0.42
-minimax-m2 302 0.35 0.86 0.51 191 0.29 0.85 0.56
-deepseek-v3.2-reasoner 277 0.43 0.80 0.37 169 0.40 0.80 0.40
-All models 1440 0.40 0.78 0.38 712 0.37 0.80 0.43
+Model #Tasks First pos. Last pos. Span #Tasks First pos. Last pos. Span
+Claude 4.5 314 0.34 0.75 0.41 101 0.30∗∗∗ 0.78∗∗ 0.48∗∗
+Gemini 3 Pro 235 0.53 0.67 0.14 73 0.55∗∗∗ 0.70∗∗∗ 0.15∗∗∗
+Kimi K2-T 309 0.40 0.82 0.42 178 0.40∗∗∗ 0.82∗∗∗ 0.42∗∗∗
+MiniMax M2 302 0.35 0.86 0.51 191 0.29∗∗∗ 0.85∗∗ 0.56∗∗
+DeepSeek v3.2-R 277 0.43 0.80 0.37 169 0.40∗∗∗ 0.80∗∗∗ 0.40∗∗∗
+All models 1440 0.40 0.78 0.38 712 0.37∗∗∗ 0.80∗∗∗ 0.43∗∗∗
 Notes.
-Values are macro-over-tasks means.
-“First test-writing position” and “Last test-writing position” are the normalized positions of the first and last test-writing events relative to total_steps.
-“Test-writing span” is the normalized distance between the first and last test-writing positions.
+Macro-over-tasks means over tasks with tests.
+Asterisks mark significant resolved-vs.-unresolved differences within a model (two-sided Mann–Whitney U: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$).
 Results.
@@
 Models differ in when they start writing tests.
-For example, gemini-3-pro-preview starts later (0.53–0.55), while minimax-m2 and claude-opus-4-5 start earlier (0.29–0.35).
+For example, Gemini 3 Pro starts later (0.53–0.55), while MiniMax M2 and Claude Opus 4.5 start earlier (0.29–0.35).
 Most models finish test writing late in the task (last position around 0.75–0.86).
 Models also differ in how spread out test writing is.
-gemini-3-pro-preview has a short span (0.14–0.15).
-minimax-m2 has a wider span (0.51–0.56), and claude-opus-4-5 is also relatively wide (0.41–0.48).
-kimi-k2-thinking is almost identical between resolved and unresolved tasks (0.40–0.82; span 0.42).
+Gemini 3 Pro has a short span (0.14–0.15).
+MiniMax M2 has a wider span (0.51–0.56), and Claude Opus 4.5 is also relatively wide (0.41–0.48).
+Kimi K2 Thinking is almost identical between resolved and unresolved tasks (0.40–0.82; span 0.42).
 Overall, unresolved tasks have a slightly larger average span than resolved tasks (0.43 vs.
 0.38).
-RQ1.2 Test-Writing Timing: Key Pattern Test writing typically finishes late, but its start time and span are mainly model-dependent; unresolved tasks are only slightly more spread out (0.43 vs. 0.38).
-3.3. RQ1.3 Execution: How Intensively Are Agent-Written Tests Executed, and With What Process Outcomes?
+RQ1.2 Test-Writing Timing: Key Pattern Test writing typically finishes late, while its start time and span are mainly model-dependent; unresolved tasks are only slightly more spread out.
+$3.$ RQ1.3 Execution: How Intensively Are Agent-Written Tests Executed, and With What Process Outcomes?
 Goal and measurements.
@@
 Resolved Unresolved
-Model #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%) #Tasks w/ tests Mean ExecCount Mean ExecPerTest Mean FailRate (%)
-claude-opus-4-5 314 4.87 1.50 11.97 101 6.27 1.68 11.14
-gemini-3-pro-preview 235 2.71 1.51 8.53 73 2.79 1.40 7.08
-kimi-k2-thinking 309 5.39 1.62 24.95 178 6.54 1.76 21.05
-minimax-m2 302 7.19 1.55 24.11 191 9.70 2.09 24.10
-deepseek-v3.2-reasoner 277 3.74 1.11 27.37 169 4.66 1.32 29.55
-All models 1440 4.89 1.46 19.68 712 6.52 1.70 21.05
+Model #Tasks w/ tests Exec Count Exec PerTest FailRate (%) #Tasks w/ tests Exec Count Exec PerTest FailRate (%)
+Claude 4.5 314 4.87 1.50 11.97 101 6.27∗∗∗ 1.68 11.14
+Gemini 3 Pro 235 2.71 1.51 8.53 73 2.79 1.40 7.08
+Kimi K2-T 309 5.39 1.62 24.95 178 6.54 1.76 21.05
+MiniMax M2 302 7.19 1.55 24.11 191 9.70∗∗∗ 2.09∗∗ 24.10
+DeepSeek v3.2-R 277 3.74 1.11 27.37 169 4.66 1.32 29.55
+All models 1440 4.89 1.46 19.68 712 6.52∗∗∗ 1.70∗∗ 21.05
 Notes.
-#Tasks: tasks with test writing.
-Mean ExecCount: test executions per task.
-Mean ExecPerTest: executions per agent-written test artifact.
-Mean FailRate: % executions with non-zero return codes (macro-over-tasks).
+#Tasks: tasks with test writing; all other values are macro-over-tasks means.
+Asterisks mark significant resolved-vs.-unresolved differences within a model (two-sided Mann–Whitney U: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$).
 Results.
@@
 1.46).
+Mann–Whitney U tests show that these aggregate resolved-vs.-unresolved differences are statistically significant for ExecCount and ExecPerTest, but not for FailRate.
+At the model level, significant differences appear mainly for Claude Opus 4.5 (higher ExecCount in unresolved tasks) and MiniMax M2 (higher ExecCount and ExecPerTest in unresolved tasks).
 FailRate is slightly higher for unresolved tasks (21.05% vs.
-19.68%).
+19.68%), but this aggregate difference is not statistically significant.
 Models differ strongly in execution intensity.
-gemini-3-pro-preview runs tests the least (ExecCount $\approx$ 2.7–2.8).
-minimax-m2 runs tests the most, especially for unresolved tasks (ExecCount 9.70; ExecPerTest 2.09).
+Gemini 3 Pro runs tests the least (ExecCount $\approx$ 2.7–2.8).
+MiniMax M2 runs tests the most, especially for unresolved tasks (ExecCount 9.70; ExecPerTest 2.09).
 FailRate also varies by model.
-claude-opus-4-5 and gemini-3-pro-preview have lower FailRate (about 7–12%), while deepseek-v3.2-reasoner, kimi-k2-thinking, and minimax-m2 are higher (about 21–30%).
-RQ1.3 Test Execution: Key Pattern For tasks that write tests, unresolved tasks run tests more often and rerun them more per written test artifact (ExecCount 6.52 vs.
-4.89; ExecPerTest 1.70 vs.
-1.46), while process-level execution failures vary mainly by model (FailRate $\sim$7%–30%).
+Claude Opus 4.5 and Gemini 3 Pro have lower FailRate (about 7–12%), while DeepSeek v3.2 Reasoner, Kimi K2 Thinking, and MiniMax M2 are higher (about 21–30%).
+RQ1.3 Test Execution: Key Pattern For tasks that write tests, unresolved runs execute them more intensively, and these aggregate differences are statistically significant, while process-level execution failures vary mainly by model.
 3.4. Summary of RQ1.
-RQ1 provides a descriptive baseline of agent-written testing behaviors in a high-autonomy setting.
-Test writing is common for most models, but extremely rare for gpt-5.2 (Table 1).
-For tasks that write tests, test writing usually ends late in the task, while the start position and the writing span vary mainly by model (Table 2).
-Unresolved tasks are slightly more spread out in test writing on average (span 0.43 vs.
-0.38).
-Finally, among tasks that write tests, unresolved tasks execute tests more often and rerun them more per written test artifact (Table 3).
-Process-level execution failures also vary by model (about 7–30% FailRate).
-RQ1, however, only describes when and how often tests appear and run, not what feedback they provide.
-RQ2 therefore analyzes the feedback encoded in agent-written tests.
-4. RQ2: What Feedback Signals Do Agent-Written Tests Provide, and What Types of Assertions Do They Use?
+RQ1 shows that agent-written testing is better understood as a model-dependent process behavior than as a simple marker of eventual success.
+This, in turn, raises a more informative question for RQ2: when these tests do appear, what kind of feedback do they actually provide?
+4. RQ2: What Feedback Signals Do Agent-Written Tests Provide?
 Motivation.
@@
 Experiment Design.
-RQ2 conditions on tasks that write at least one test artifact and reports descriptive summaries of two aspects of test feedback:
-  • Signal counts (RQ2.1): how many feedback statements appear in agent-written tests, split into assertions vs. value-revealing prints.
+RQ2 conditions on tasks that write at least one test artifact and reports descriptive summaries of three aspects of test feedback:
+  • Signal counts (RQ2.1): assertion vs. value-revealing print counts in agent-written tests.
   • Assertion types (RQ2.2): what types of assertions appear in agent-written tests, using a four-type categorization.
+  • Print types (RQ2.3): what value-revealing prints typically inspect, using a three-type categorization.
+RQ2.1 stays at the task level and measures how much feedback agent-written tests emit overall.
+RQ2.2 and RQ2.3 then move to the statement level, categorizing assertion and value-revealing print statements separately.
 4.1. RQ2.1 Task-level feedback signal amount: How much feedback do agent-written tests encode?
@@
 Conditioning on tasks that contain agent-written test artifact, we quantify how many feedback statements are encoded in those artifacts.
-We distinguish two signal types: (i) verification signals ($A$), i.e., assert statements that specify explicit checks, and (ii) observational signals ($P$), i.e., value-revealing print statements that expose runtime values or computed expressions.
+We distinguish two signal types: (i) verification signals (AA), i.e., assert statements that specify explicit checks, and (ii) observational signals ($P$), i.e., value-revealing print statements that expose runtime values or computed expressions.
 To ensure $P$ (prints) reflects observational feedback, we exclude pure-literal prints that emit only fixed strings (e.g., print("here")) and count only prints that expose runtime values, expressions, or execution results (e.g., print(obj.attr)).
@@
 Table 4. Task-level feedback signal amount encoded in agent-written tests
-Resolved Unresolved Model #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) claude-opus-4-5 314 5.16 25.00 30.16 101 5.36 25.61 30.97 gemini-3-pro-preview 235 1.45 4.34 5.79 73 1.62 5.04 6.66 kimi-k2-thinking 309 2.86 20.72 23.57 178 3.51 24.03 27.54 minimax-m2 302 7.37 34.06 41.43 191 4.66 43.09 47.76 deepseek-v3.2-reasoner 277 3.51 16.43 19.94 169 3.31 20.95 24.27
+Resolved Unresolved Model #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) #Tasks w/ tests Assertions per task ($\bar{N}^{A}$) Prints per task ($\bar{N}^{P}$) Total signals per task ($\bar{N}^{\mathrm{total}}$) Claude 4.5 314 5.16 25.00 30.16 101 5.36 25.61 30.97 Gemini 3 Pro 235 1.45 4.34 5.79 73 1.62 5.04 6.66 Kimi K2-T 309 2.86 20.72 23.57 178 3.51 24.03 27.54 MiniMax M2 302 7.37 34.06 41.43 191 4.66 43.09 47.76 DeepSeek v3.2-R 277 3.51 16.43 19.94 169 3.31 20.95 24.27
 Note. Macro-over-tasks means computed over tasks with tests. Assertions count assert statements; prints count value-revealing prints. $\bar{N}^{\mathrm{total}}=\bar{N}^{A}+\bar{N}^{P}$.
-Figure 2. Composition of feedback signals in agent-written tests across models. Value-revealing prints dominate over assertions for all models.
+Figure 2. Composition of feedback signals in agent-written tests across models.
 Results.
-Table 4 shows that, when agent-written tests are present, they can contain a substantial number of feedback statements per task (e.g., 19.94–47.76 total signals for several models).
+Table 4 shows that, when agent-written tests are present, they can contain a substantial number of feedback statements per task.
 As shown in Figure 2, feedback is predominantly observational: for every model, value-revealing prints exceed assertions in the macro-average counts per task.
-Models differ markedly in overall signal volume.
-For example, minimax-m2 encodes the largest total signal counts (41.43–47.76), whereas gemini-3-pro-preview encodes far fewer (5.79–6.66), with other models in between.
-Across models, unresolved tasks tend to show slightly higher total signal counts, driven mainly by more value-revealing prints (e.g., deepseek-v3.2-reasoner: +4.33 total; minimax-m2: +6.33 total).
-Assertion counts are comparatively stable; notably, minimax-m2 has fewer assertions but more prints on unresolved tasks.
-RQ2.1 Test Signal Amount: Key Takeaway When agents write tests, those tests often emit many runtime feedback statements per task, but the feedback is mostly observational: value-revealing prints consistently outnumber assertions across all models.
-Signal volume varies widely by model (e.g., minimax-m2 high vs.
-gemini-3-pro-preview low), and unresolved tasks show slightly more total signals mainly due to additional prints, while assertion counts remain relatively stable.
+Signal volume also varies markedly by model, from the lower totals of Gemini 3 Pro to the much higher totals of MiniMax M2.
+Across models, unresolved tasks tend to show slightly higher total signal counts, driven mainly by more value-revealing prints, while assertion counts remain comparatively stable.
+RQ2.1 Test Signal Amount: Key Takeaway When agents write tests, the feedback is mostly observational: value-revealing prints consistently outnumber assertions, although total signal volume still varies markedly by model.
 4.2. RQ2.2 Assertion categorization: What kinds of verification do assertions encode?
@@
   • C4 Exact checks. The assertion checks an exact value or deep structural equality. Example: assert output == expected_output.
-To identify and categorize assertions, we implement a rule-based classifier over Python ASTs and map each extracted assertion to exactly one category.
+To identify and categorize assertions, we implement a rule-based classifier over Python ASTs and map each assertion to exactly one category.
 The classifier covers both native assert statements (e.g., assert a == b) and framework-provided assertion calls (e.g., self.assertEqual(a, b) in unittest).
 Concretely, for each test artifact, we parse the code into an AST and extract assertion events from: (i) native assert <expr> statements, and (ii) calls to framework assertion APIs.
-Some assert statements contain multiple checks in one line, combined with boolean operators (e.g., assert a > 0 and b == 1).
+Some assert statements contain multiple checks in one line with boolean operators (e.g., assert a > 0 and b == 1).
 In this example, a > 0 is a constraint check (C3) and b == 1 is an exact check (C4).
-For such compound assert statements, we decompose the expression into atomic checks and assign a single category by taking the highest category present under the ordering from C1 to C4, because the most specific check in the statement best reflects what the assertion is trying to enforce.
-Thus, assert a > 0 and b == 1 is labeled as C4.
+For such compound assert statements, we decompose the expression into atomic checks and assign a single category by taking the highest category under the ordering from C1 to C4, because the most specific check in the statement best reflects what the assertion is trying to enforce.
+Thus, assert a > 0 and b == 1 is labeled C4.
 Table 5. Assertion category distribution by model. Counts and percentages are computed over all assertion statements written by each model.
-Model #Assertions C1 Sanity C2 Property C3 Relational C4 Exact # % # % # % # % claude-opus-4-5 2160 351 16.25% 807 37.36% 93 4.31% 909 42.08% gemini-3-pro-preview 458 76 16.59% 154 33.62% 36 7.86% 192 41.92% kimi-k2-thinking 1508 225 14.92% 622 41.25% 45 2.98% 616 40.85% minimax-m2 3117 618 19.83% 1291 41.42% 132 4.23% 1076 34.52% deepseek-v3.2-reasoner 1531 285 18.62% 537 35.08% 52 3.40% 657 42.91%
+Model #Assertions C1 Sanity C2 Property C3 Relational C4 Exact # % # % # % # % Claude 4.5 2160 351 16.25% 807 37.36% 93 4.31% 909 42.08% Gemini 3 Pro 458 76 16.59% 154 33.62% 36 7.86% 192 41.92% Kimi K2-T 1508 225 14.92% 622 41.25% 45 2.98% 616 40.85% MiniMax M2 3117 618 19.83% 1291 41.42% 132 4.23% 1076 34.52% DeepSeek v3.2-R 1531 285 18.62% 537 35.08% 52 3.40% 657 42.91%
 Note.
@@
 Across all five models, most assertions fall into C2 Property and C4 Exact, while C3 Relational remains consistently uncommon.
-For four models (claude-opus-4-5, gemini-3-pro-preview, kimi-k2-thinking, and deepseek-v3.2-reasoner), C4 Exact accounts for roughly 41–43% of assertions (40.85–42.91%), and C2 Property accounts for roughly 34–41% (33.62–41.25%).
-minimax-m2 follows the same overall shape but allocates a smaller share to C4 Exact (34.52%) and larger shares to C1 Sanity (19.83%) and C2 Property (41.42%).
-Across all models, C3 Relational is rare (2.98–7.86%), with the highest proportion in gemini-3-pro-preview (7.86%).
-Overall, the consistent scarcity of C3 may have several practical explanations.
-First, relational or approximate checks can be more delicate to specify and maintain than local property checks or direct equality checks.
-Second, such checks may be less common in everyday unit-test patterns that models imitate during code generation.
-Third, for many SWE-bench issues, agents may find it more straightforward to write either local property checks (C2) or exact expected outputs (C4) once they have a candidate fix.
+For four models (Claude Opus 4.5, Gemini 3 Pro, Kimi K2 Thinking, and DeepSeek v3.2 Reasoner), C4 Exact accounts for roughly 41–43% of assertions (40.85–42.91%), and C2 Property accounts for roughly 34–41% (33.62–41.25%).
+MiniMax M2 follows the same overall shape but allocates a smaller share to C4 Exact (34.52%) and larger shares to C1 Sanity (19.83%) and C2 Property (41.42%).
+Across all models, C3 Relational is rare (2.98–7.86%), with the highest proportion in Gemini 3 Pro (7.86%).
+This scarcity suggests that agents more often fall back on local property checks (C2) or exact expected outputs (C4), while relational or approximate constraints may be harder to specify and less common in the test patterns models imitate.
 We treat these distributions as descriptive of which assertion forms appear in agent-written tests, not as evidence of correctness or impact on task resolution.
-RQ2.2 Assertion Categorization: Key Takeaway Across models, assertions in agent-written tests are dominated by two forms: checks of local properties (e.g., checking that a required attribute is present) and checks against exact expected values.
-Assertions that express relational or range-style constraints are consistently uncommon (only a small single-digit fraction of assertions across models).
-minimax-m2 has a lower share of exact-value checks and a higher share of sanity-guard and local-property checks than the other models, but the overall pattern is similar across models.
-4.3. Summary of RQ2
-RQ2 shifts from when agents test (RQ1) to what feedback their agent-written tests produce at execution time.
-First, when tests are present, they often emit many feedback statements per task, but the feedback is dominated by value-revealing prints rather than assertions: across all models, print signals consistently outnumber assert checks, and total signal volume varies widely by model (RQ2.1).
-Second, when agents do write assertions, models show broadly similar mixes of assertion forms: most assertions either check local properties (e.g., the presence or validity of an attribute or field) or check exact expected values, while relational or range-style constraints are rare (RQ2.2).
-Taken together, RQ2 characterizes agent-written tests primarily as an observational feedback channel: most test feedback comes from value-revealing prints.
+RQ2.2 Assertion Categorization: Key Takeaway Across models, assertions are dominated by property checks and exact-value checks, whereas relational or range-style constraints remain uncommon.
+4.3. RQ2.3 Print categorization: What do value-revealing prints typically inspect?
+Goal and measurements.
+Since value-revealing prints substantially outnumber assertions, we further inspect what these prints expose in practice.
+This helps clarify whether agent-written tests are primarily used to inspect values, inspect structural summaries, or surface execution status, thereby refining our interpretation of tests as observational debugging tools.
+We group value-revealing prints into three categories:
+  • P1 Value / content inspection.
+The print inspects a concrete runtime value or content, such as a returned output, an intermediate result, an object field, or part of a generated string.
+This category is used when the print is meant to reveal what the program produced, rather than a structural summary or an error/status signal; e.g., print(add(1, 2)).
+  • P2 Structural summary inspection.
+The print reports an aggregate structural summary, such as length, size, shape, number of items, or emptiness.
+This category is reserved for prints that summarize how much data is present or what structure it has, rather than showing the content itself; e.g., print(len(items)).
+  • P3 Exception / execution-status signals.
+The print surfaces an exception, an error message, or a coarse execution-status indicator, such as a success/failure flag.
+This category is used when the print primarily indicates what happened during execution, rather than the program’s computed content; e.g., except Exception as exc: print(exc).
+We implement a rule-based classifier over Python ASTs.
+For each parseable test artifact, we extract value-revealing print(...) calls, exclude pure-literal or formatting-only prints, and assign each print to one of the three categories above using deterministic category rules.
+Table 6. Value-revealing print category distribution by model.
+Model #Prints P1 Value/Content P2 Structural P3 Exception/Status # % # % # % Claude 4.5 7919 6136 77.48% 274 3.46% 1509 19.06% Gemini 3 Pro 1352 942 69.67% 72 5.33% 338 25.00% Kimi K2-T 8909 6556 73.59% 584 6.56% 1769 19.86% MiniMax M2 15685 11003 70.15% 921 5.87% 3761 23.98% DeepSeek v3.2-R 7495 5709 76.17% 280 3.74% 1506 20.09%
+Results.
+Table 6 shows a stable pattern across models.
+P1 Value / content inspection dominates for every model (69.67–77.48%), indicating that most prints are used to inspect concrete outputs, intermediate values, or object contents.
+P3 Exception / execution-status signals form the second-largest category (19.06–25.00%), whereas P2 structural-summary inspection remains comparatively uncommon (3.46–6.56%).
+Overall, value-revealing prints are used mainly to inspect runtime values and coarse execution outcomes rather than encode strong pass/fail criteria, reinforcing our interpretation of agent-written tests as observational debugging tools.
+RQ2.3 Print Categorization: Key Takeaway Across models, most value-revealing prints are used for value/content inspection, with exception/execution-status signals as a distant second.
+Structural summaries are much less common, reinforcing that these prints function primarily as observational debugging probes rather than strong correctness oracles.
+$4.$ Summary of RQ2
+RQ2 shows that agent-written tests function primarily as an observational feedback channel: value-revealing prints dominate, and the assertions that do appear are concentrated in local-property and exact-value checks.
 This naturally raises the next question: do these agent-written tests meaningfully affect task resolution?
-We address this in RQ3.
-5. RQ3: Do Agent-Written Tests Truly Affect Task Resolution?
+5. RQ3: How Does Prompting Test Writing Change Observed Outcomes and Costs?
 Motivation.
@@
 For example, GPT-5.2 almost never writes new test artifacts (3/500 tasks, 0.6%), yet it still resolves 71.8% of tasks.
-In contrast, claude-opus-4.5 writes at least one new test artifact in about 83% of tasks, but its resolution rate is only 2.6 percentage points higher (74.4%).
+In contrast, Claude Opus 4.5 writes at least one new test artifact in about 83% of tasks, but its resolution rate is only 2.6 percentage points higher (74.4%).
 RQ2 further shows that, when tests are written, most feedback comes from value-revealing prints rather than assert-based checks.
-These findings raise a direct question: Do writing tests truly affect task resolution outcomes, and at what cost?
+These findings raise a direct question: when prompts change whether agents write tests, how do the observed task outcomes and costs change in this setting?
 Experiment Design.
 RQ3 answers two questions:
-  • RQ3.1 (Outcome impact): If we encourage or discourage agents to write tests, how do task resolution outcomes change?
-  • RQ3.2 (Efficiency impact): If we encourage or discourage agents to write tests, how do API calls and token usage change?
+  • RQ3.1 (Outcome changes): If we encourage or discourage agents to write tests, how do observed task resolution outcomes change?
+  • RQ3.2 (Efficiency changes): If we encourage or discourage agents to write tests, how do API calls and token usage change?
 Model selection.
-To isolate the effect of agent-written tests, we design two complementary intervention experiments: (i) encouraging agents to write tests, and (ii) discouraging agents from writing new test files.
+To probe how prompting test writing changes observed behavior under the same scaffold, we design two complementary intervention experiments: (i) encouraging agents to write tests, and (ii) discouraging agents from writing new test files.
 We choose models for each setup based on their baseline test-writing rate observed in RQ1 (Table 1), defined as the fraction of tasks where the agent writes test artifacts.
 For the encourage test writing setup, we focus on low test-writing models and medium test-writing models, so there is meaningful headroom for increasing test creation.
+We do not include already test-heavy models in this setup, because their baseline test-writing rates leave little room for a meaningful upward shift.
 Specifically, we include GPT-5.2 (0.6%), an extreme low test-writing model in RQ1 with near-zero test creation.
-We also include gemini-3-pro-preview (61.1%), a medium test-writing model whose baseline test creation is already substantial but still leaves room for further increase.
+We also include Gemini 3 Pro (61.1%), a medium test-writing model whose baseline test creation is already substantial but still leaves room for further increase.
 For the discourage test writing setup, we start from high test-writing models that write tests in the vast majority of tasks in RQ1: four models show consistently high test-writing rates (83.0%–98.6%; Table 1).
-Due to budget constraints, we select two representatives from this group: kimi-k2-thinking (97.4%) and deepseek-v3.2-reasoner (89.2%).
-Concretely, we use two prompt variants:
-  • Encourage writing tests: for gpt-5.2 and gemini-3-pro-preview, we append an instruction to write at least one runnable new test file (a file whose name starts with test_ or ends with _test.py), separate from the repository’s existing tests.
-  • Discourage writing tests: for kimi-k2-thinking and deepseek-v3.2-reasoner, we (i) remove the sentence “Test edge cases to ensure your fix is robust” and (ii) append an instruction to not write any new test files or scripts; robustness and edge cases should be handled using reasoning and code inspection only.
-By comparing the performance of each agent under these revised prompts against their original performance, we isolate the specific impact of test writing.
-Consequently, the observed changes in behaviors or outcomes can be directly attributed to the agent-written tests.
-5.1. RQ3.1 Does encouraging or discouraging test writing change task resolution?
+Due to budget constraints, we select two representatives from this group: Kimi K2 Thinking (97.4%) and DeepSeek v3.2 Reasoner (89.2%).
+Concretely, we start from the original mini-SWE-agent prompt and make small, targeted prompt edits to create two variants:
+  • Encourage writing tests: for GPT-5.2 and Gemini 3 Pro, we append prompt instructions to write at least one runnable new test file (a file whose name starts with test_ or ends with _test.py), separate from the repository’s existing tests.
+  • Discourage writing tests: for Kimi K2 Thinking and DeepSeek v3.2 Reasoner, we remove the default testing-related prompt cue from mini-SWE-agent and append prompt instructions not to write any new test files or scripts.
+The exact texts of the original prompt and both revised variants are included in the experiment_prompts folder.
+[^2] Specifically, the files are mini_swe_agent_original_prompt.yaml, encourage_write_tests_prompt.yaml, and discourage_write_tests_prompt.yaml.
+We compare each revised prompt against its baseline to measure how changing the prompt is accompanied by changes in test writing, outcomes, and costs.
+5.1. RQ3.1 How does encouraging or discouraging test writing change observed task outcomes?
 Goal and measurements.
-To answer whether encouraging or discouraging test writing changes task resolution, for each model, we compare each task under two conditions: the baseline run (under standard mini-SWE-agent prompt) and the intervention run (under our revised prompt).
-Specifically, we record two features: whether the run creates at least one new test artifact (No test vs.
-Has test) and whether the patch successfully resolve the issue (Fail vs.
-Success).
-Then, we analyze how these two features change from the baseline run to the intervention run.
-This respectively results in four possible transition groups for both test writing (No $\rightarrowNo test$, No $\rightarrowHas test$, Has $\rightarrowNo test$, Has $\rightarrowHas test$) and task outcome (Fail $\rightarrow$ Success, Success $\rightarrow$ Fail, Stable Success, and Stable Fail).
-To visualize the relationship between these shifts, we represent the results using a transition matrix.
-In this matrix, the rows represent the change in test-writing behavior, while the columns represent the change in task outcomes.
-This structure allows us to pinpoint the exact impact of our intervention.
-For instance, the intersection of No test $\rightarrow$ Has test and Fail $\rightarrow$ Success represents the instances where encouraging test-writing directly led to a breakthrough in resolving the task.
+[…]
 Table 7. Test-writing status flips and outcome transitions
@@
 Our prompt interventions substantially change whether models write test artifacts, but these shifts rarely translate into outcome changes.
-As shown in Table 6, the encourage test writing prompt flips test status for the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview: 64.4% and 37.0% of tasks transition from No test to Has test, respectively.
-Conversely, the discourage test writing prompt removes tests at scale for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, moving 68.4% and 75.2% of tasks from Has test to No test.
+As shown in Table 7, the encourage test writing prompt flips test status for the low test-writing model GPT-5.2 and the medium test-writing model Gemini 3 Pro: 64.4% and 37.0% of tasks transition from No test to Has test, respectively.
+Conversely, the discourage test writing prompt removes tests at scale for the high test-writing models Kimi K2 Thinking and DeepSeek v3.2 Reasoner, moving 68.4% and 75.2% of tasks from Has test to No test.
 Despite these large test-status shifts, resolution outcomes are largely stable.
@@
 Table 7 further indicates that even when test status flips, success rates change only slightly.
-For example, for deepseek-v3.2-reasoner, discouraging test writing removes tests in 376 tasks but yields a net decrease of only 20 resolved tasks, a small change relative to the behavioral shift.
+The exact McNemar tests reinforce this descriptive picture: none of the four models shows a statistically significant baseline-vs.-intervention outcome shift (all $p>0.05$).
+For example, for DeepSeek v3.2 Reasoner, discouraging test writing removes tests in 376 tasks but yields a net decrease of only 20 resolved tasks, a small change relative to the behavioral shift.
 Overall, changing how often a model writes test artifacts appears to be a weak lever for shifting task outcomes in this setting.
+5.1.1. Exploring Issue Types for Which Tests May Help
+[…]
+Results.
+We find 8 issues that appear in both categories: astropy-13236; django-11790, django-13401, django-13512, and django-14493; pylint-7080; scikit-learn-14629; and sympy-21612.
+We manually read the problem statements of these 8 issues and used GPT-5.4 to help summarize recurring themes.
+This exploratory pass suggests three recurring properties: they are mostly small-to-moderate correctness bugs, often involve explicit reproduction conditions or edge-case triggers, and frequently require checking precise expected behavior or semantics.
+These observations are necessarily limited.
+SWE-bench Verified contains 500 tasks in total, the benchmark does not provide an official issue taxonomy, and our cross-category overlap contains only 8 instances.
+We therefore present this subsection as an exploratory step beyond the main intervention result, intended to surface plausible test-sensitive candidates rather than to establish definitive issue categories.
+A broader taxonomy-based analysis is left to future work.
 RQ3.1: Outcomes vs.
-Test-Status Changes Discouraging test writing removes agent-written tests at scale (at least 68% of tasks for kimi-k2-thinking and deepseek-v3.2-reasoner).
-Yet these large behavioral shifts rarely translate into outcome shifts: across models, 83.2% of tasks keep the same final resolution result after the intervention.
-Even when tests are induced, success does not increase correspondingly, suggesting that simply writing more test artifacts provides little leverage for fixing the underlying issue.
-Conversely, suppressing tests for hundreds of tasks does not cause outcomes to collapse; for deepseek-v3.2-reasoner, removing tests in 376 tasks yields only 20 fewer successes.
+Test-Status Changes Prompt interventions can flip test-writing status at scale, yet most tasks keep the same final outcome.
+Even large induced or suppressed shifts in test writing therefore appear to be a weak lever on task resolution in this setting.
 5.2. RQ3.2 How do API calls and token usage change?
@@
 We further analyze the following three metrics based on the trajectories generated in RQ3.1: (i) average API calls per task, (ii) average input tokens per task, and (iii) average output tokens per task.
-Table 7. API calls and token usage under baseline vs. encourage-tests / discourage-tests conditions.
-Model Condition Tasks resolved Avg API Calls Avg Input Tokens Avg Output Tokens gpt-5.2 Baseline 359 (71.8%) 19.76 242,855 24,550 Encourage writing tests 359 (71.8%) 20.84 264,762 29,415 Change +0 (+0.0%) +1.08 (+5.5%) +21,907 (+9.0%) +4,866 (+19.8%) gemini-3-pro-preview Baseline 371 (74.2%) 40.33 666,096 11,114 Encourage writing tests 366 (73.2%) 39.21 641,307 10,943 Change -5 (-1.0%) -1.11 (-2.8%) -24,789 (-3.7%) -171 (-1.5%) kimi-k2-thinking Baseline 317 (63.4%) 46.82 668,449 14,895 Discourage writing tests 304 (60.8%) 30.25 340,689 8,468 Change -13 (-2.6%) -16.57 (-35.4%) -327,760 (-49.0%) -6,427 (-43.1%) deepseek-v3.2-reasoner Baseline 300 (60.0%) 46.40 637,297 52,120 Discourage writing tests 291 (58.2%) 35.06 427,780 44,823 Change -9 (-1.8%) -11.35 (-24.5%) -209,518 (-32.9%) -7,297 (-14.0%)
+For each model and metric, we compare the intervention run against the baseline run on the same 500 tasks.
+To quantify whether the observed efficiency changes are directionally robust, we additionally compute two-sided Wilcoxon signed-rank tests and 95% paired bootstrap confidence intervals for the task-level difference (Condition - Baseline).
+Table 8. API calls and token usage under baseline vs. test-encouragement / test-discouragement conditions.
+Model Condition Tasks resolved Avg API Calls Avg Input Tokens Avg Output Tokens GPT-5.2 Baseline 359 (71.8%) 19.76 242,855 24,550 Encourage tests 359 (71.8%) 20.84 264,762 29,415 Change -+0 (+0.0%) -+1.08 (+5.5%)∗∗∗ -+21,907 (+9.0%)∗∗∗ -+4,866 (+19.8%)∗∗∗ Gemini 3 Pro Baseline 371 (74.2%) 40.33 666,096 11,114 Encourage tests 366 (73.2%) 39.21 641,307 10,943 Change -5 (-1.0%) -1.11 (-2.8%) -24,789 (-3.7%) -171 (-1.5%) Kimi K2-T Baseline 317 (63.4%) 46.82 668,449 14,895 Discourage tests 304 (60.8%) 30.25 340,689 8,468 Change -13 (-2.6%) -16.57 (-35.4%)∗∗∗ -327,760 (-49.0%)∗∗∗ -6,427 (-43.1%)∗∗∗ DeepSeek v3.2-R Baseline 300 (60.0%) 46.40 637,297 52,120 Discourage tests 291 (58.2%) 35.06 427,780 44,823 Change -9 (-1.8%) -11.35 (-24.5%)∗∗∗ -209,518 (-32.9%)∗∗∗ -7,297 (-14.0%)∗∗∗
 Note.
 Changes are computed as (Condition - Baseline).
-Encourage writing tests is applied to gpt-5.2 and gemini-3-pro-preview; Discourage writing tests is applied to kimi-k2-thinking and deepseek-v3.2-reasoner.
+Encourage tests is applied to GPT-5.2 and Gemini 3 Pro; Discourage tests is applied to Kimi K2-T and DeepSeek v3.2-R.
+$^{***}p<0.001$ by paired Wilcoxon signed-rank test; paired bootstrap CIs are summarized in the text.
 Results.
-Table 7 reports task outcomes (Tasks resolved) and efficiency metrics (API calls and tokens) under the baseline and the two intervention conditions.
-Overall, the interventions have only marginal impact on resolution rates, but they can noticeably reshape efficiency—in ways that depend on each model’s baseline test-writing propensity.
-For the encourage test writing setup, the effects differ between the low test-writing model gpt-5.2 and the medium test-writing model gemini-3-pro-preview.
-For gpt-5.2, encouraging test writing increases overhead: API calls rise by 5.5% and output tokens by 19.8%, while the resolution rate remains unchanged from baseline.
-In the medium test-writing model gemini-3-pro-preview, we observe very small shifts in overall usage (e.g., $1.5\%$ completion tokens).
-A likely reason is that this model already writes tests in about 61.6% of baseline tasks, leaving limited headroom for the prompt to flip tasks from No test to Has test; accordingly, the net change in testing-related interaction is small on average.
-Moreover, because the model is nondeterministic, even when a task contains tests in both baseline and encouraged runs, the amount of testing (and related interaction) can differ slightly between the paired runs; these small within-task fluctuations average out, keeping aggregate cost differences modest and making a small decrease reasonable.
-The most striking efficiency shifts appear in the discourage test writing setup for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner.
-Discouraging test writing yields large reductions in resource consumption, with input tokens dropping by 49.0% and 32.9%, respectively.
-For kimi-k2-thinking, average API calls fall by 35.4%, cutting interaction workload by over a third.
-Notably, these efficiency gains come with only small decreases in success rates (2.6% for kimi-k2-thinking and 1.8% for deepseek-v3.2-reasoner), suggesting that a substantial portion of the baseline interaction budget spent on test writing for these models has limited marginal benefit for task resolution in this setting.
-RQ3.2: Cost changes are larger than outcome changes Changing whether agents write tests has little effect on task resolution, but it can noticeably change efficiency.
-More tests increases overhead for low test-writing model models without improving outcomes: for gpt-5.2, API calls rise by 5.5% and output tokens by 19.8% while the resolution rate stays unchanged.
-In contrast, suppressing tests yields large efficiency gains for high test-writing models: input tokens drop by 49.0% (kimi-k2-thinking) and 32.9% (deepseek-v3.2-reasoner), and kimi-k2-thinking reduces API calls by 35.4%.
-These savings come with only small success drops (2.6% and 1.8%), implying that most test-writing overhead in these models delivers low marginal utility for producing a successful final patch.
+Table 8 shows that the interventions have only marginal impact on resolution rates, but can noticeably reshape efficiency.
+Under encourage test writing, the low test-writing model GPT-5.2 incurs higher overhead (+5.5% API calls; +19.8% output tokens) without any gain in resolution.
+The paired analyses support this increase across all three metrics: API calls rise by +1.08 per task (95% CI [0.39, 1.75], Wilcoxon $p<0.001$), input tokens by +21,907 ([3,954, 39,598], $p<0.001$), and output tokens by +4,866 ([3,041, 6,704], $p<0.001$).
+In contrast, the medium test-writing model Gemini 3 Pro changes little overall, and none of its paired differences reaches conventional significance across API calls, input tokens, or output tokens (all 95% CIs cross zero).
+The most striking shifts appear under discourage test writing for the high test-writing models Kimi K2 Thinking and DeepSeek v3.2 Reasoner: input tokens drop by 49.0% and 32.9%, respectively, and Kimi K2 Thinking also reduces API calls by 35.4%.
+These reductions are directionally robust in the paired analyses: for both models, all three mean differences are negative, all 95% bootstrap CIs exclude zero, and all Wilcoxon tests give $p<0.001$.
+These paired results therefore strengthen the interpretation that suppressing test writing yields substantial efficiency savings in these two high test-writing models, whereas encouraging test writing produces model-dependent and much less stable cost effects.
+RQ3.2: Cost changes are larger than outcome changes Changing whether agents write tests has much larger effects on efficiency than on task resolution.
+Paired Wilcoxon tests and bootstrap confidence intervals support the large discourage-setting savings, while encourage-setting effects are robust for GPT-5.2 but small and statistically unclear for Gemini 3 Pro.
 Summary of RQ3.
-Overall, more agent-written tests do not mean more solves in this high-autonomy setting.
-When we push models to write more tests, test-writing status flips at scale (e.g., 64.4% for gpt-5.2), yet task outcomes remain largely unchanged: on average, 83.2% of tasks keep the same success/fail result after the intervention.
-At the same time, more tests can be expensive: for the low test-writing model gpt-5.2, inducing test writing increases overhead (+5.5% API calls; +19.8% output tokens) without any gain in resolution.
-Conversely, fewer tests can be much cheaper with only small outcome losses: for the high test-writing models kimi-k2-thinking and deepseek-v3.2-reasoner, suppressing test writing cuts input tokens by 49.0% and 32.9%, respectively, while success drops remain modest (2.6% and 1.8%).
-Taken together, varying the amount of agent-written tests strongly reshapes resource usage but has limited leverage on whether the final patch resolves the issue, suggesting that much of the test-writing effort provides low marginal utility for task resolution.
+Overall, varying the amount of agent-written tests strongly reshapes resource usage but has limited leverage on whether the final patch resolves the issue.
+More tests do not mean more solves in this high-autonomy setting, but they can impose substantial interaction overhead.
 6. Discussion and Future Work
 6.1. Implications
-Our study indicates that the testing behaviors conducted by current LLMs, without clear guidance, do not show significant improvement in agent-based software issue resolution.
-It suggests two main implications for the research and practice in designing and utilizing LLM software agents.
-On the one hand, our results reveal inherent limitations in how existing LLMs leverage testing for issue resolution when explicit guidance is absent.
-This highlights the importance of instructing agents not only on whether tests should be written, but also on how they should be designed and, more critically, how to effectively interpret and utilize the feedback generated by test execution.
-On the other hand, given that current LLMs often fail to extract meaningful value from test writing, practitioners may consider adopting a more conservative approach to agent-generated tests.
-Such an approach could involve more sophisticated prompting strategies or guardrail mechanisms that help agents determine when test writing is truly necessary.
-Additionally, code agents could benefit from incorporating a cost–benefit monitoring framework that tracks the overhead associated with test-related activities, such as test writing, execution, and failure analysis, to assess their actual contribution to patch refinement.
+Our results suggest three practical implications.
+For practitioners, the goal should not be to simply make agents write more tests, but to make testing more targeted and budget-aware.
+One lightweight approach is to package reusable testing behaviors as focused Claude Code subagents/skills (Anthropic, 2026b, a), such as generating one minimal regression test, strengthening a weak oracle by converting prints into assertions, or running only the smallest relevant test slice.
+It is also useful to log test-writing, test-running, and failure-analysis costs separately so teams can see when extra test-related interaction stops adding value.
+For researchers, the main implication is to study testing as a process intervention, not only as a final success-rate difference.
+Beyond reporting outcome transitions, future studies should trace where validation budget is spent across test writing, test execution, failure inspection, and patch revision.
+In practice, this can be done with observability tooling such as LangSmith tracing (LangChain, 2026), making it easier to ask when testing helps, what feedback is useful, and when cheaper validation behaviors may be enough.
+For benchmark users and maintainers, our study points to a narrower measurement concern.
+In its February 23, 2026 article “Why SWE-bench Verified no longer measures frontier coding capabilities,” OpenAI argued that SWE-bench Verified has become harder to interpret because contamination and residual benchmark-design issues weaken the meaning of final scores (OpenAI, 2026).
+We do not evaluate those benchmark-level claims directly, nor do we test contamination or training-data effects.
+Our evidence instead supports a more limited point: even when final resolution changes little, agents can still differ substantially in how often they write tests, how much validation budget they spend, and how their software-engineering behavior unfolds.
+This suggests that final solve rates on SWE-bench Verified are too coarse as a standalone measure of agent behavior and are better interpreted alongside process-sensitive metrics.
 6.2. Future work
-Our findings motivate two future-work topics on agent-written testing in lightly scaffolded, high-autonomy setups.
+Two future directions follow from our results.
 Evaluating on-the-fly test quality in a non-stationary code state.
 Traditional test-quality metrics (e.g., coverage, mutation score, fault revelation) assume a fixed snapshot of the system under test and reproducible executions (Yu et al., 2023; Ryan et al., 2024; Shin et al., 2024; Bhatia et al., 2024; Yang et al., 2024b; Molinelli et al., 2025; Harman et al., 2025).
-In agentic development, the code state is non-stationary: tests are written and run against intermediate repository versions that may later be overwritten, and the final patch may not preserve the exact state a test originally validated.
-This complicates attribution (which version/function a test exercised), reproducibility, and the use of conventional quality pipelines.
-Future work should develop instrumentation for execution-time evaluation—e.g., capturing runnable snapshots (code, environment, commands) at test time—and define metrics that remain meaningful for transient intermediate artifacts.
-Self-evolving test-generation strategies without over-constraining exploration.
-Our results highlight a tension between structure and autonomy: fully self-directed testing can increase cost with limited outcome gains, while a heavy, fixed workflow can narrow the search space and preclude better strategies.
-The goal is to produce higher-value tests by deciding when to test, what to verify (oracle design), and how to allocate budget across reasoning, editing, and validation, while preserving task-specific flexibility.
-A promising direction is self-evolving (Robeyns et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Hu et al., 2025) testing policies: instead of a static hand-written prompt, allow the agent to revise its own testing prompt/policy from environment feedback and failure modes, adapting to task context and model capabilities (Gao et al., 2025a).
-Future work can formalize this as closed-loop optimization under cost and safety constraints, and compare human-specified versus self-adapted strategies under matched budgets and controlled scaffolds.
+In agentic development, tests are written and run against intermediate repository versions that may later be overwritten, complicating attribution and reproducibility.
+Future work should therefore develop execution-time instrumentation and metrics that remain meaningful for transient intermediate artifacts.
+Self-evolving test-generation strategies.
+A promising direction is self-evolving (Robeyns et al., 2025; Zhang et al., 2025a; Xia et al., 2025; Hu et al., 2025) testing policies, in which the agent revises its own testing prompt or policy from environment feedback and failure modes rather than following a static hand-written prompt (Gao et al., 2025a).
+Future work can formalize this as closed-loop optimization under cost and safety constraints, comparing human-specified and self-adapted strategies under matched budgets and controlled scaffolds.
 6.3. Threats to Validity
@@
 Moreover, success–failure comparisons can reflect differences in task difficulty or interaction length (e.g., debugging duration), not only testing.
-We mitigate these concerns by treating observational results as descriptive, by using prompt-only interventions that change test creation while keeping the agent setup fixed, and by reporting task-level outcome transitions (not only aggregate resolution deltas) to show where outcomes change versus remain stable.
+We mitigate these concerns by treating observational results as descriptive, by using prompt-only interventions that keep the agent setup fixed, and by reporting task-level outcome transitions rather than only aggregate resolution deltas.
 External validity.
 Our findings are based on SWE-bench under a light scaffold and a specific set of models and providers; absolute magnitudes may differ under other benchmarks, programming languages, toolchains (e.g., enforced CI), or future model versions.
-To support transfer, we focus on patterns that may recur in similar agent settings (e.g., large cross-model differences in testing style, observation-dominant feedback, and limited outcome sensitivity under sizable shifts in test creation), and we provide precise measurement definitions and intervention prompts to facilitate replication under alternative setups.
+To support transfer, we focus on patterns that may recur in similar agent settings (e.g., large cross-model differences in testing style, observation-dominant feedback, and limited outcome sensitivity under sizable shifts in test creation), and we provide precise measurement definitions and intervention prompts for replication.
 Data construction validity.
@@
 These procedures can miss unconventional test artifacts, project-specific helpers, or edge-case syntax patterns.
-We reduce construction error by using AST parsing (rather than surface regex heuristics), conservative rules for counting value-bearing prints, and fully deterministic, reproducible extraction and categorization; nevertheless, results should be interpreted with respect to these explicit definitions.
+We reduce construction error through AST parsing, conservative rules for counting value-bearing prints, and deterministic extraction and categorization; nevertheless, results should be interpreted with respect to these definitions.
 7. Related Work
@@
 In contrast, our study focuses on agent-written tests that emerge dynamically during high-autonomy, multi-step resolution of real-world GitHub issues, where the codebase and candidate patches evolve over time.
-We treat test writing and execution as an emergent process behavior, and characterize (i) whether/when/how intensively agents test, (ii) what signals tests encode (assertions vs.
-observational prints), and (iii) how these behaviors relate to resolution outcomes.
+We treat test writing and execution as an emergent process behavior and characterize both the signals these tests encode and how those behaviors relate to resolution outcomes.
 Trajectory Analysis of Software Agents.
@@
 Overall, existing trajectory analyses emphasize action sequences, outcome separation, and error categorization, but rarely examine whether and how agents autonomously decide to test.
-Our work addresses this gap by characterizing the emergent testing behaviors observed in execution trajectories and analyzing what feedback agent-written tests actually provide during issue resolution.
+Our work addresses this gap by characterizing emergent testing behaviors and the feedback agent-written tests provide during issue resolution.
 8. Conclusion
 This paper revisited the common intuition that ”testing helps” for LLM-based software agents in a high-autonomy setting where writing and running tests is not specified in the prompt.
-Using SWE-bench Verified trajectories produced under a light agent scaffold, we separated (i) whether/when/how agents write and execute tests, (ii) what feedback those tests encode, and (iii) whether changing test-writing instructions actually shifts task outcomes and efficiency.
-RQ1 established a descriptive baseline of emergent testing behaviors.
-Across models, test writing propensity differs dramatically (from near-universal test writing to almost none), while within the same model, resolved and unresolved tasks show broadly similar test-writing rates.
-When tests are present, unresolved tasks tend to spread test writing across a slightly larger portion of the trajectory and execute tests more often, with higher re-execution intensity; process-level execution failures, however, vary mainly by model rather than outcome.
-These results highlight that in a high-autonomy setting, ”testing” is a model-specific process behavior and is not tightly coupled to eventual success.
-RQ2 then showed that the content of agent-written tests is largely observational.
-Across all models, value-revealing prints consistently outnumber assert-based checks, and overall signal volume varies widely by model.
-When assertions are used, their forms are broadly similar across models: most are local property checks or exact-value checks, while relational or range-style constraints are consistently rare.
-Taken together, agent-written tests function primarily as a runtime probing interface rather than a systematic verification mechanism.
-Finally, RQ3 provided controlled evidence about impact.
-Prompt-only interventions flip test-writing status at scale—inducing tests for the low test writing model and suppressing test writing for high test writing models—yet task outcomes are mostly unchanged (the majority of tasks preserve the same success/fail result).
-In contrast, efficiency effects can be substantial: for gpt-5.2—the near-zero test-writing model—encouraging test writing increases overhead without improving resolution, whereas for test-heavy models (e.g., kimi-k2-thinking and deepseek-v3.2-reasoner) discouraging test writing can sharply reduce API calls and token usage with only modest decreases in success.
-Overall, more agent-written tests do not mean more solves: varying test-writing effort primarily reshapes resource usage and interaction patterns, with limited leverage on whether the final patch resolves the issue.
-Overall, our findings provide an empirical answer to the question: Help or Habit?
-In high-autonomy software issue resolution, agent-written tests often behave more like a reproduced software development lifecycle routine than a dependable source of help: models that write many tests are not consistently more successful, and prompt interventions that induce or suppress test writing rarely change whether the issue is resolved.
-What tests do change is the process footprint—API calls, token usage, and interaction steps—indicating that test writing frequently reflects how an agent chooses to work, not whether it can reliably validate the patch.
+Across our three research questions, agent-written testing is better understood as a model-dependent process style than as a dependable driver of success: test-writing propensity varies sharply across models, test feedback is dominated by value-revealing prints rather than assertions, and prompt-only changes in test-writing cues usually have little effect on observed task outcomes even when they substantially change efficiency.
+Overall, these findings suggest that agent-written tests often behave more like a habitual software-development routine than a dependable source of validation in this setting.
+More agent-written tests do not mean more solves; what they more reliably change is the process footprint—API calls, token usage, and interaction patterns.
+Improving the value of testing for code agents may therefore require better oracles and more actionable validation signals, rather than simply inducing agents to write more tests.
 9. Data Availability
-To support transparency and reproducibility, we publicly release a replication package containing the datasets, raw trajectories, and analysis scripts used in this study.
-The package is available on Figshare: https://figshare.com/s/8fa6412b86bf54fcd27e?file=61474897.
+We provide an anonymous Zenodo replication package containing the datasets, raw trajectories, prompt files, and analysis scripts used in this study. DOI: Phttps://doi.org/10.5281/zenodo.19251470